### v1

# A new model of spontaneous scalarization of black holes induced by curvature and matter

Abstract
We propose a new model of black hole spontaneous scalarization that combines a scalar-Gauss-Bonnet interaction with a non-minimal coupling to a U(1) gauge field (a dark photon or an electromagnetic field). This construction generalizes earlier single-coupling setups and allows both curvature-induced and matter-induced scalarization within one framework, which allows us to overcome the limitations of each mechanism alone. We focus on charged, spherically symmetric black holes and demonstrate that our model substantially expands the range of black hole masses and charges that permit scalar hair. Negative Gauss-Bonnet couplings, previously associated only with near-extremal charges or rapidly spinning black holes, now trigger scalarization for much broader charge intervals. We develop a numerical procedure to solve the field equations, and investigate the various properties of these black holes. This results in new branches emerging at distinct mass thresholds, a behavior not seen in the pure Einstein-scalar-Gauss–Bonnet or Einstein-scalar-Gauss-Bonnet-Ricci models. The scalar charge depends sensitively on the coupling parameters and on the $U(1)$ charge. Our analysis also shows that these black holes have larger entropy than their Reissner–Nordström counterparts and can become overcharged, surpassing the usual extremal limit of charge-to-mass ratio. Analyzing the scalar charge behavior suggests that adding matter-coupling appears to stabilize solutions that were previously prone to higher-order instabilities in pure Gauss–Bonnet models with quadratic coupling and broadens the range of possible configurations, making this model a promising candidate for further studies in strong gravity.
I Introduction
Gravitational-wave observations have transformed black hole research and opened a new era of gravitational astronomy. Since 2015 the LIGO-Virgo-KAGRA collaboration has recorded more than 200 binary black hole mergers and one binary neutron star merger, beginning with the landmark GW150914 [1]. These signals reveal highly nonlinear spacetime dynamics and offer direct measurements of the strong-gravity regime. Complementary data from the Event Horizon Telescope enrich this picture, showing the shadow of the supermassive black hole in M87* [2] and the central black hole of our galaxy Sagittarius A* [3]. Future images will tighten constraints on horizon-scale geometry. These data allow tests of General Relativity (GR) and the Kerr hypothesis [4] by comparing predicted and observed waveform models and shadow profiles. Future gravitational-wave detectors such as LISA [5, 6, 7], Einstein Telescope [8, 9] and Cosmic Explorer with improved sensitivity will enlarge the catalog of detections, and extend the mass and redshift ranges. They might probe signatures of alternative gravitational theories which were not observed before.
Black holes (BH) in GR are characterized by only three quantities: mass, spin and electric charge. No-hair theorems reinforce this claim and exclude extra scalar charges for some models when gravity couples to matter in the minimal way [10, 11]. Their proofs extend to many models [12], so the metric is like GR and the scalar field is trivial. However, if one assumption of the no-hair theorems is violated, a black hole can carry scalar hair. Early attempts led to pathological, unstable solutions (e.g. the Bocharova-Bronnikov-Melnikov-Bekenstein (BBMB) solution [13, 14]), but later analyses found regular solutions when the coupling meets certain conditions. Identifying which couplings lead to stable, regular solutions that match observations remains essential.
GR has been extraordinarily successful in weak-field regimes, but its predictions in the strong-field regime remain largely untested. This experimental gap, now being addressed by gravitational-wave observations comes alongside theoretical puzzles that motivate extending GR. For example, phenomena like galaxy rotation curves (attributed to dark matter) and the accelerated cosmic expansion (dark energy) are not explained by GR with known fields. In addition, we cannot reconcile GR with quantum mechanics. These issues have prompted modifications of Einstein’s theory, such as adding higher-curvature terms to the action or introducing new degrees of freedom. Such extensions can emerge from fundamental physics, such as some attempts to unify gravity with other interactions naturally. A broad range of modified gravity models has been developed (see these review papers [15, 16, 17, 18, 19]), and upcoming strong-field tests offer a way to distinguish these alternative theories from GR in regimes where deviations may become pronounced.
One of the principal frameworks for modifying gravity is the class of scalar-tensor theories that introduce a dynamical scalar field alongside the tensor field of GR. The prototypical example, Brans-Dicke theory, was formulated to incorporate Mach’s principle by allowing the gravitational constant to vary [20]. Modern scalar-tensor models are constructed to reduce to GR in appropriate limits (ensuring consistency with solar-system experiments) while offering richer behavior in large-scale structure or in other extreme conditions. In particular, they can provide an effective explanation for cosmic acceleration without a cosmological constant, all while obeying current local gravity constraints. The most general form of these theories is Horndeski theories [21, 22, 23, 24, 25]. They are designed so that their field equations remain second-order, avoiding ghost instabilities. No-hair theorems were developed for some classes of scalar-tensor theories. Hawking proved it for Brans-Dicke theories [10], then [16] extend the proof to $f(R)$ theories. Hui and Nicolis established no-hair theorems for most interesting classes of Horndeski theories [26]. Certain scalar-tensor theories allow the black hole to carry a nontrivial scalar field configuration. In other words they admit "hairy" black holes solutions that evade no-hair theorems. As proved in these theorems scalar hair is forbidden in many scalar-tensor models unless specific conditions are met, as when the scalar is coupled to additional invariants like Gauss-Bonnet (GB) term [27]. The possibility of scalar hair on black holes marks a significant departure from standard GR and motivates further empirical tests in the strong gravity regime.
Higher-curvature corrections involving the GB invariant (a particular combination of squared curvature terms) arise naturally in several fundamental contexts. Notably, a GB term appears in the low-energy effective action of heterotic string theory [28, 29, 30, 31, 32] and in higher-dimensional generalizations of GR (Lovelock gravity) [33, 34]. In four dimensions the GB term by itself is a topological invariant and does not affect the field equations, but if it is coupled to a scalar field one obtains a nontrivial extension known as Einstein-scalar-Gauss-Bonnet (EsGB) gravity. EsGB models have second-order field equations and remain free of ghost instabilities, forming a subclass of Horndeski theories [23]. A well-studied example is Einstein-dilaton-Gauss-Bonnet (EdGB), which employs an exponential coupling $e^{-\alpha\phi}$ motivated by string theory . Hairy black hole solutions in this model were first found perturbatively in [35] and then numerically in [36]. After that [27] showed that shift symmetric coupling of the scalar field to GB invariant allow to get hairy solutions. Recently [37, 38] discovered a new violation of no-hair theorems in EsGB theories for a large class of coupling functions including a quadratic coupling . EsGB theories provide a promising arena to explore strong-gravity phenomena and to confront theoretical predictions with observational data.
Another important category of alternative gravity involves a scalar field non-minimally coupled to the electromagnetic field. In such Einstein-Maxwell-Scalar (EMS) models, the scalar is coupled non-minimally to the Maxwell invariant through a coupling function $g(\phi)$. These models have very well motivated theoretical origins: Einstein-Maxwell-dilaton (EMD) theory for instance, emerges from Kaluza-Klein compactification and also appears as a low-energy limit of string theory, as well as in Supergravity frameworks. Another strong historical feature of such theories is that electromagnetic fields can source the scalar, leading to black hole solutions different from the usual Reissner-Nordström (RN) solution of GR. A well-known example is the Gibbons-Maeda-Garfinkle-Horowitz-Strominger dilaton black hole in EMD gravity [39, 40]. This theory does not admit the standard RN (bald) solution with scalar hair, any static charged BH must carry a "hair". The dilaton coupling thereby alters the black hole’s charge-mass relation; BH can attain charge-to-mass ratios exceeding the extremal limit of RN black holes (they are overcharged). Models with scalar-Maxwell couplings, often inspired by high energy theory, thus provide a rich framework for how additional fields might influence electromagnetic interactions in strong gravity, offering distinctive black holes signatures that can be probed observationally.
One of the most recent and actively studied mechanisms that allow compact objects to develop scalar hair only under some extreme conditions is spontaneous scalarization (see [41] for a review). Damour and Esposito-Farèse first uncovered the mechanism for neutron stars [42]. In their model a massless scalar field is coupled to the trace of matter stress-energy tensor. When the stellar compactness reaches a critical value of the coupling, it drives a tachyonic instability. The scalar grows until nonlinear terms stop the process, leaving the star with a non-trivial configuration of scalar field "scalar hair". The scalar field is dormant in weak fields and still agrees with solar-system tests but emerges for dense configurations. Later it was shown [43, 44] that spontaneous scalarization can apply to vacuum spacetimes (black holes) if the scalar couples to curvature or other matter invariants.
Two broad classes of models realize this phenomenon: one is the curvature-induced scalarization through Gauss-Bonnet coupling [43, 44], the other is matter-induced scalarization through couplings to field such electromagnetism [45, 46]. In both cases, a nonminimal coupling of a scalar field $\phi$ to an additional invariant (geometric invariant such as the GB term, the Ricci scalar or the Chern-Simons term [47, 48, 49], or matter invariant such as the electromagnetic tensor $F_{\mu\nu}F^{\mu\nu}$) triggers a tachyonic instability when the invariant is sufficiently large, causing the GR solution (with $\phi=0$) to become unstable, inducing a new ground state with $\phi\neq 0$. This provides a natural way to circumvent no-hair theorems by introducing scalar hair only beyond a certain threshold of curvature or field strength. In the curvature-induced case, represented by EsGB theories, the scalar field is coupled to the Gauss–Bonnet invariant ${\cal G}$ (a quadratic curvature term). A representative action is
$S_{\rm EsGB}=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\Big{[}R-2\nabla_{\mu}\phi\nabla^{\mu}\phi+f(\phi){\cal G}\Big{]}~{}$
with $f(\phi)$ a coupling function. The Schwarzschild black hole is a solution with $\phi=0$; however, if $f(\phi)$ is chosen such that $f^{\prime\prime}(0)\,{\cal G}>0$ induces a tachyonic mass for $\phi$, then above some coupling strength or below a certain mass scale the Schwarzschild solution becomes unstable. Indeed, it was shown that a static vacuum black hole can spontaneously scalarize when its curvature (characterized by $M^{-2}$ for a black hole of mass $M$) exceeds a critical value [44, 43]. The endpoint is a “hairy” black hole with a nontrivial $\phi(r)$ profile. A key feature of these Gauss–Bonnet models is that they preserve the GR solutions for weak curvature (large $M$), activating scalar hair only in the high-curvature regime. This behavior is motivated by the idea that deviations from GR appear only in extreme gravity environments (as might be expected from quantum gravity or high-energy effective corrections), thereby satisfying observational constraints in the weak-field regime, while providing new physics for compact objects. This is why curvature-induced scalarization of black holes has attracted a lot of interest [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. Another attractive aspect is that the Gauss–Bonnet term is a natural curvature correction arising in effective descriptions (it appears in low-energy expansions of string theory), so EsGB models introduce scalar hair through a well-motivated extension of the gravitational action.
Likewise, spontaneous scalarization can be induced by couplings to matter fields. An important example is the EMS class of models [45, 46, 69, 70], where the scalar field couples to the electromagnetic field invariant $F_{\mu\nu}F^{\mu\nu}$. In these models a charged Reissner–Nordström black hole (which in electrovacuum has $\phi=0$) can become unstable if the electric field is strong enough, leading to a new branch of scalarized charged black holes. For instance, with an action term of the form $f(\phi)F_{\mu\nu}F^{\mu\nu}$, the Maxwell field can trigger a tachyonic growth of $\phi$ when the charge $Q$ of the black hole exceeds a certain threshold. This matter-induced scalarization is analogous to the Gauss–Bonnet case: the scalar hair develops only when the electric field (or charge-to-mass ratio $Q/M$) is sufficiently large, ensuring that minimally coupled GR solutions remain stable for weakly charged black holes. Such EMS models have been extensively studied [71, 72, 73, 73, 74, 75, 61, 76, 77, 78, 79, 80, 81, 82, 83] as a way to generate hairy black holes that are smooth deformations of the Reissner–Nordström solution, again obeying the principle that hair appears only in regimes where Maxwell field is above some threshold.
There have been some attempts to study charged black holes with Gauss-Bonnet-induced scalarization, without any direct scalar-Maxwell coupling [84, 85, 86]. These studies found that the inclusion of charge can enrich the scalarization behavior. For example, the Gauss–Bonnet term admits two types of scalarization (depending on the sign of the coupling) $GB^{+}$ or $GB^{-}$, and while a spherical symmetric neutral black hole only scalarizes in the positive-coupling case, adding charge allows scalarization with negative coupling ($GB^{-}$), but only near extremality. The electromagnetic field in those models influences the curvature-induced mechanism indirectly by modifying the black hole spacetime, but it does not act as an independent source of scalarization because $\phi$ does not couple to $F_{\mu\nu}F^{\mu\nu}$. Thus, the curvature-driven instability remains the primary route to hair in these models.
A crucial challenge for the viability of EsGB scalarization is the stability of the scalarized solutions. It has become evident that many black holes with Gauss-Bonnet-induced hair suffer from perturbative instabilities [87]. In particular, for the simple choice of a quadratic coupling function $f(\phi)\propto\phi^{2}$, all the resulting scalarized black hole solutions are unstable under radial (spherically symmetric) perturbations [88, 89]. This means that although the hair can form, it does not represent a stable end state for evolution, small perturbations will drive the solution away (often back to the GR state). More sophisticated coupling functions (for example, exponential [43, 88] or higher-order forms [90, 89]) can produce branches of solutions that are radially stable. However, even those “stable” branches have been shown to be vulnerable to non-radial (angular) instabilities. In fact, modes with higher angular momentum can become unstable on hairy solutions that are stable to monopole perturbations. For instance, in theories of EsGB with Ricci coupled quadratically to $\phi$ refereed as (EsGBR) [91, 92], a Quadrupole ($l=2$) instability was found by [93], which can be extended to EsGB theories. Another instability was reported in EsGB and in EsGBR, called gradient instability, where the sound speed of angular propagation becomes imaginary for even perturbations in the case of large harmonic index $l$ [94]. The same authors later mentioned a mistake in their previous calculation, and found that the instability occurs for only part of the parameter space [95]. Many references reported a hyperbolicity loss in EsGB and EsGBR scalarization models and analyzed its threshold using numerical simulations [96, 97, 98, 99, 100, 101].
These instability issues imply that achieving a truly stable scalarized black hole in Gauss–Bonnet models is nontrivial. The existence of such instabilities motivates searching for other alternative models that can stabilize the solutions.
Motivated by these developments, in this work we propose a new model of black hole spontaneous scalarization induced by both curvature and matter, aiming to capture both curvature-induced and matter-induced scalarization within a single framework. The model introduces a simultaneous coupling of the scalar field to the Gauss–Bonnet invariant and to $U(1)$ gauge field (a dark photon or an electromagnetic field), incorporating the two mechanisms within one framework. Our aim is to leverage the two sources of scalarization to overcome some of the difficulties faced by the pure EsGB case.
By including the matter-coupling directly, the scalar field has an additional channel to become tachyonic, which may widen the parameter space for stable scalarized solutions or provide new branches of hair that are absent in single-coupling models. Intuitively, the curvature coupling ensures that strong-gravity environments can trigger the scalar, while the matter coupling adds an extra contribution that may cure instabilities and modify the properties of the resulting black holes. This unified model is a logical next step because it reduces to the known cases in the appropriate limits (recovering pure Gauss-Bonnet scalarization when charge is zero, and pure Maxwell-induced scalarization when the Gauss-Bonnet coupling is turned off). At the same time, it offers the prospect of richer phenomenology by allowing a broader scalarization threshold for different masses and, potentially, more robust hairy black hole solutions. Another motivation, comes from the fact that this model can be seen as a subset of Maxwell-Horndeski theories, whose linear stability has already been studied [102]. This will allow us to check easily if the model suffers from the same instabilities reported previously. We should add that by analogy to similar cases this model might be seen as a low energy limit of string actions [35].
Finally, let us clarify the interpretation of the $U(1)$ gauge field. We treat this field and the associated charge in a broad sense without fixing its physical interpretation. The standard view is to consider this $U(1)$ field as the usual electromagnetic field. If the black hole is electrically charged, astrophysical arguments imply that net electric charge stays negligible when the black hole interacts with ordinary plasmas and light charged particles. In that setting the charge-to-mass ratio is very small, or the theory serves as a toy model to study the impact of additional interactions with EsGB models of scalarization.
Another possibility is a magnetically charged black hole  [^1] We consider only electrically charged BHs in this paper, but our calculations can be extended to magnetic BHs and dyonic BHs, motivated by hypothetical magnetic monopoles if magnetic monopoles exist, either primordial or formed through other early-universe processes or motivated by GUT-based arguments. In that case, black holes might preserve large magnetic charge over long timescales because there is no abundant light magnetic counterpart to neutralize it. This has attracted much interest recently [103, 104, 105, 106, 107, 108], especially after the paper of Maldacena [109].
Alternatively, one may view the $U(1)$ field as living in a hidden sector decoupled from the Standard model (dark photon) [110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128], with black holes carrying a dark charge that evades standard discharge arguments because the charge-to-mass ratio of dark particles can be much larger or smaller. We can identify the gauge field with a small fractional charge (minicharged dark matter) that might endow black holes with a net hidden charge [129]. Various other references studied black holes involving dark photons [130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141].
Consequently, the model remains agnostic about how the black hole acquired its charge, whether it is electrical, magnetic, or hidden, and accommodates several theoretical and phenomenological contexts under a single framework.
From a theoretical point of view, RN black holes are interesting because they admit an extremal limit, so they can also serve as a toy model for understanding rotating solutions. Among four-dimensional Einstein-Maxwell solutions [142], RN black holes have a unique property: in the extremal limit, they admit a static, regular multi-black-hole configuration (the Majumdar-Papapetrou solution [143, 144] that remains regular on and outside the event horizon). In that scenario, gravitational attraction and electromagnetic repulsion cancel exactly at every point. This condition often occurs in supersymmetric contexts. The extremal RN solution is the only four-dimensional Einstein-Maxwell black hole with Killing spinors when viewed as the bosonic sector of $\mathcal{N}=2$ Supergravity [145, 146].
In the following sections, we detail the construction of this model with two sources and explore its black hole solutions and their properties. Section II introduces the model and field equations, then analyzes the onset of scalarization, Section III presents numerical solutions of the scalarized black holes their physical properties like the scalar charge and thermodynamic properties, and Section IV discusses conclusions and possible future directions.
In this study, we adopt the convention $4\pi G=1=4\pi\epsilon_{0}$ for ease of calculation. The spacetime signature is (,+,+,+)(-,+,+,+)( - , + , + , + ). Our analysis is limited to spherically symmetric solutions, meaning that all metric and matter fields are functions of the radial coordinate only. To simplify our notation, any function initially introduced as dependent on the radial coordinate, like $X(r)$, will henceforth be referred to simply as $X$, keeping in mind its radial dependence. Derivatives are denoted as $X^{\prime}\equiv\frac{dX}{dr}$ for radial derivatives and $X_{,\phi}\equiv\frac{dX}{d\phi}$ for derivatives with respect to the scalar field $\phi$.
II Setup : Theoretical framework
II.1 Description of the model
We consider a family of theories [^2] These theories can be considered as a generalization of Einstein-Maxwell-scalar-Gauss-Bonnet theories, and a subclass of Maxwell-Horndeski theories. where a real scalar field couples non-minimally both to the Gauss-Bonnet invariant and to a $U(1)$ gauge field (e.g. an electromagnetic field).
The action is :
$S=\frac{1}{16\pi}\int d^{4}x\,\sqrt{-g}\Big{[}R-2\nabla_{\mu}\phi\nabla^{\mu}\phi+f(\phi)\mathcal{G}-g(\phi)F_{\mu\nu}F^{\mu\nu}\Big{]}\ ,$ (1)
with $R$ the Ricci scalar and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ where $A_{\mu}$ is a $U(1)$ gauge field, $F_{\mu\nu}$ is its strength tensor (that can be identified as a Maxwell tensor), $\phi$ a scalar field non-minimally coupled to gravity and $U(1)$ gauge field through the functions $f(\phi)$ and $g(\phi)$ respectively, and $\mathcal{G}$ is the Gauss-Bonnet invariant which is defined as follows:
$\mathcal{G}=R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}\ .$ (2)
Variation of the action with respect to the metric and the different fields leads to the following equations:
  • Einstein equations:
    $G_{\mu\nu}=T_{\mu\nu}^{\rm(\phi)}+T_{\mu\nu}^{\rm(EM)}$ (3)
    with
    $T_{\mu\nu}^{\rm(\phi)}=2(\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}\partial_{\alpha}\phi\partial^{\alpha}\phi\ -2P_{\mu\gamma\nu\alpha}\nabla^{\alpha}\nabla^{\gamma}f(\phi))\ ,$ (4)
    $T_{\mu\nu}^{\rm(EM)}=2g(\phi)(F_{\mu\alpha}F_{\nu}^{\ \alpha}-\frac{1}{4}g_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta})\ ,$ (5)
    where
    $P_{\alpha\beta\mu\nu}=-\frac{1}{4}\varepsilon_{\alpha\beta\rho\sigma}R^{\rho\sigma\gamma\delta}\varepsilon_{\mu\nu\gamma\delta}$ $=R_{\alpha\beta\mu\nu}+g_{\alpha\nu}R_{\beta\mu}-g_{\alpha\mu}R_{\beta\nu}+g_{\beta\mu}R_{\alpha\nu}-g_{\beta\nu}R_{\alpha\mu}$ (6)
    $+\frac{1}{2}\left(g_{\alpha\mu}g_{\beta\nu}-g_{\alpha\nu}g_{\beta\mu}\right)R\ ,$
    and $\varepsilon_{\alpha\beta\rho\sigma}$ is the Levi-Civita tensor.
  • Klein-Gordon equation:
    $\Box\phi+\frac{1}{4}f_{,\phi}(\phi)\mathcal{G}-\frac{1}{4}g_{,\phi}(\phi)F_{\mu\nu}F^{\mu\nu}=0$ (7)
  • Maxwell’s equations:
    $g(\phi)\nabla_{\nu}F^{\mu\nu}+g_{,\phi}(\phi)F^{\mu\nu}\nabla_{\nu}\phi=0$ (8)
II.2 Field equations and ansatz
We focus on spherically symmetric models. A convenient metric ansatz is
$ds^{2}=-N(r)\sigma(r)^{2}dt^{2}+\frac{dr^{2}}{N(r)}+r^{2}\big{(}d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}\big{)}\ ,~{}~{}{\rm with}~{}~{}N(r)\equiv 1-\frac{2\,m(r)}{r},$ (9)
where $m(r)$ is the Misner-Sharp mass function [147] and $\sigma(r)$ a second metric function. Spherical symmetry, in the absence of a magnetic charge [^3] A magnetic charge would also be compatible with spherical symmetry, but it shall not be considered here - see [70] for dyonic charged BHs in the context of EMS theories., imposes an electrostatic 4-vector potential, $A=V(r)\,dt$, and a scalar field that depends only on $r$, $\phi=\phi(r)$.
Substituting into the field equations leads to a set of three coupled differential equations for the radial functions $N$, $\sigma$, and $\phi$, with an additional constraint. This system of equations can be put in a form suitable for numerical integration:
$N^{\prime}=F_{1}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),\qquad$ $\sigma^{\prime}=F_{2}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),$ (10)
$\phi^{\prime\prime}=F_{3}(N,\sigma,\phi,\phi^{\prime}),\quad$ $(\frac{V^{\prime}g(\phi)r^{2}}{\sigma})^{\prime}=0.$
In more detail, these equations can be written as
$N^{\prime}$ $=\frac{-r^{2}g(\phi)\left(N\left(\phi^{\prime 2}\left(r^{2}-4(N-1)f_{,\phi\phi}(\phi)\right)-4(N-1)\phi^{\prime\prime}f_{,\phi}(\phi)+1\right)-1\right)-Q^{2}}{r^{2}g(\phi)\left(2(1-3N)\phi^{\prime}f_{,\phi}(\phi)+r\right)},$ (11)
$\sigma^{\prime}$ $=\frac{\sigma\left(\phi^{\prime 2}\left(r^{2}-8(N-1)f_{,\phi\phi}(\phi)\right)-8(N-1)\phi^{\prime\prime}f_{,\phi}(\phi)\right)}{2(1-3N)\phi^{\prime}f_{,\phi}(\phi)+r},$
$V^{\prime}$ $=-\frac{Q\sigma}{g(\phi)r^{2}},$
$\phi^{\prime\prime}$ $=\frac{W}{Z},$
The electrostatic potential $V$ gives rise to a first integral, simplifying the remaining equations. After integration of the last equation of (10), the integration constant $Q$ will be interpreted as the electric charge.
The detailed expressions for $W$ and $Z$ are reported in Appendix A.
II.3 Boundary conditions and physical quantities of interest
To solve the set of four coupled ordinary differential equations (11), one must implement the appropriate boundary conditions. We impose regularity at the horizon, $r=r_{H}$, the field equations can be approximated by a power series expansion in $r-r_{H}$ as
$N=N_{0}+N_{1}(r-r_{H})+\cdots\ \ ,\qquad\sigma=\sigma_{0}+\sigma_{1}(r-r_{H})+\cdots\ ,$ (12)
$\phi=\phi_{0}+\phi_{1}(r-r_{H})+\cdots\ ,\qquad\ \ V=V_{1}(r-r_{H})+\cdots\ ,$
where the coefficients are given by
$N_{0}$ $=0,$ $N_{1}$ $=\frac{r_{H}^{2}g(\phi_{0})-Q^{2}}{r_{H}^{2}g(\phi_{0})\left(2\phi_{1}f_{,\phi}(\phi_{0})+r_{H}\right)},$ (13)
$\sigma_{1}$ $=\frac{\sigma_{0}\phi_{1}^{2}\left(2f_{,\phi\phi}(\phi_{0})+r_{H}^{2}\right)}{2\phi_{1}f_{,\phi}(\phi_{0})+r_{H}},$ $\phi_{1}$ $=\frac{B\pm\sqrt{\Delta}}{A},$
with
$B$ $=Q^{2}r_{H}^{2}\left(g(\phi_{0})\left(4f_{,\phi}(\phi_{0})^{2}+r_{H}^{4}\right)-2r_{H}^{2}f_{,\phi}(\phi_{0})g_{,\phi}(\phi_{0})\right)+4Q^{4}g_{,\phi}(\phi_{0})^{2}-r_{H}^{8}g(\phi_{0})^{2},$
$\Delta$ $=\left(Q^{2}-r_{H}^{2}g(\phi_{0})\right)^{2}\Big{(}8Q^{2}f_{,\phi}(\phi_{0})^{3}\left(2Q^{2}f_{,\phi}(\phi_{0})-3r_{H}^{4}g_{,\phi}(\phi_{0})\right)$
$\quad+16Q^{2}r_{H}^{2}g(\phi_{0})f_{,\phi}(\phi_{0})^{2}\left(6f_{,\phi}(\phi_{0})^{2}+r_{H}^{4}\right)+r_{H}^{8}g(\phi_{0})^{2}\left(r_{H}^{4}-24f_{,\phi}(\phi_{0})^{2}\right)\Big{)},$
$A$ $=4r_{H}f_{,\phi}(\phi_{0})\left(Q^{2}r_{H}^{2}f_{,\phi}(\phi_{0})g_{,\phi}(\phi_{0})-Q^{2}g(\phi_{0})\left(4f_{,\phi}(\phi_{0})^{2}+r_{H}^{4}\right)+r_{H}^{6}g(\phi_{0})^{2}\right).$
The coefficients depend on the two parameters $\phi_{0}$ and $\sigma_{0}$, where the subscript 00 denotes functions evaluated at the horizon $r_{H}$.
The equation determining $\phi_{1}$ is obtained from the scalar field equation near the horizon and has two solutions, but we choose the only one that gives a black hole solution $\phi_{1}=\frac{B+\sqrt{\Delta}}{A}$, while the other solution $\phi_{1}=\frac{B-\sqrt{\Delta}}{A}$ does not lead to an asymptotically flat BH [36]. To obtain a physical solution we need to ensure regularity by imposing that $\Delta\geq 0$ which keeps $\phi_{1}$ real and $\phi^{\prime\prime}_{0}$ finite [36].
At spatial infinity we expand in power series of $1/r$, so the leading order expansion for far fields is
$N\approx 1-\frac{2M}{r},\qquad\sigma\approx 1-\frac{D^{2}}{2r^{2}},\qquad$ $\phi\approx\frac{D}{r},\qquad\ \ V\approx\psi_{e}+\frac{Q}{r}$ (14)
Where $M$ is the Arnowitt-Deser-Misner (ADM) mass, $D$ is the scalar charge and $\psi_{e}$ is the electrostatic potential at infinity.
We can check the regularity of the solutions and the presence of singularities in some cases by checking the divergence of the Ricci scalar $R$, and the Kretschmann scalar $K\equiv R_{\mu\nu\delta\lambda}R^{\mu\nu\delta\lambda}$,
$R=-\frac{\sigma(r)\left(-2+2N(r)+4rN^{\prime}(r)+r^{2}N^{\prime\prime}(r)\right)+r\left(3rN^{\prime}(r)\sigma^{\prime}(r)+2N(r)\left(2\sigma^{\prime}(r)+r\sigma^{\prime\prime}(r)\right)\right)}{r^{2}\sigma(r)}.$ (15)
$K$ $=\frac{4\,(N(r)-1)^{2}\,\sigma(r)^{2}+2r^{2}\sigma(r)^{2}N^{\prime}(r)^{2}+2r^{2}\!\bigl{(}\sigma(r)N^{\prime}(r)+2N(r)\sigma^{\prime}(r)\bigr{)}^{2}}{r^{4}\sigma(r)^{2}}$ (16)
$+\frac{r^{4}\!\bigl{(}3N^{\prime}(r)\sigma^{\prime}(r)+\sigma(r)N^{\prime\prime}(r)+2N(r)\sigma^{\prime\prime}(r)\bigr{)}^{2}}{r^{4}\sigma(r)^{2}}.$
Scaling symmetry
The field equations of the model exhibit a scaling symmetry [86, 44, 56], so that the solutions are invariant under the rescaling transformations
$r\to\beta r\,,\quad M\to\beta M\,,\quad\eta\to\beta^{2}\eta\,,\quad Q\;\to\;\beta\,Q,$ (17)
where $\beta>0$ is an arbitrary scaling parameter, $\eta$ is the coupling constant of the GB term, defined so that $f=\eta\,\tilde{f}$, with dimensionless $\tilde{f}$, so that $\eta$ has dimension of length square.
This symmetry and freedom in choosing the length scale implies that physical quantities of the black hole solution should be expressed in terms of dimensionless combinations that remain invariant under these transformations. To characterize the black hole solutions we introduce the reduced quantities :
$q=\frac{Q}{M}\ ,\quad\frac{\eta}{M^{2}},\quad\frac{D}{\sqrt{\eta}},\quad\frac{Q}{\sqrt{\eta}},\quad\frac{M}{\sqrt{\eta}},\quad\frac{\eta}{r_{H}^{2}}$ (18)
For numerical analysis, we exploit this symmetry to fix one parameter, typically either the horizon radius, $r_{H}=1$, or the mass, $M=1$, or the coupling parameter $\eta$ when appropriate, reducing the solution space to a family characterized by dimensionless quantities like $\eta/M^{2}$, without loss of generality.
All dimensional quantities presented in the results are normalized accordingly, either by appropriate powers of $M$ or $\sqrt{\eta}$ or $r_{H}$.
II.4 Conditions for scalarization: Scalarization threshold and coupling functions
The scalar field $\phi$ obeys the Klein-Gordon equation
$\Box\phi=\frac{-\mathcal{I}_{GB}}{4}f_{,\phi}+\frac{\mathcal{I}_{M}}{4}g_{,\phi},$ (19)
where we have renamed
$\mathcal{I}_{M}=F_{\mu\nu}F^{\mu\nu}\ ,\quad\mathcal{I}_{GB}=\mathcal{G}.$ (20)
Spontaneous scalarization occurs when the theory admits solutions with $\phi=0$. This demands
$\frac{df}{d\phi}\Big{|}_{\phi=0}=0,\quad\frac{dg}{d\phi}\Big{|}_{\phi=0}=0.$ (21)
These conditions are generally met in models with $\mathbb{Z}_{2}$ symmetry (under $\phi\rightarrow-\phi$) and allow the existence of standard electro-vacuum GR solutions.
A transition from a scalar-free black hole to one that supports a nontrivial scalar can occur if the second derivatives of $f(\phi)$ and $g(\phi)$ near $\phi=0$ create a tachyonic instability. Small-$\phi$ expansions for both coupling functions show:
$f(\phi)=f(0)+\frac{1}{2}\frac{d^{2}f}{d\phi^{2}}\Big{|}_{\phi=0}\phi^{2}+\cdots,\quad g(\phi)=g(0)+\frac{1}{2}\frac{d^{2}g}{d\phi^{2}}\Big{|}_{\phi=0}\phi^{2}+\cdots,$ (22)
The linearized the Klein-Gordon equation then becomes
$(\Box-\mu_{\rm eff}^{2})\delta\phi=0,\quad\text{where}\quad\mu_{\rm eff}^{2}=\mu_{\rm eff,f}^{2}+\mu_{\rm eff,g}^{2}$ (23)
with
$\mu_{\rm eff,f}^{2}=\frac{-1}{4}\frac{d^{2}f}{d\phi^{2}}\Big{|}_{\phi=0}\mathcal{I}_{GB},\quad\mu_{\rm eff,g}^{2}=\frac{1}{4}\frac{d^{2}g}{d\phi^{2}}\Big{|}_{\phi=0}\mathcal{I}_{M}$ (24)
and $\mu_{\rm eff,f}^{2}$ and $\mu_{\rm eff,g}^{2}$ are the effective mass squares.
A tachyonic instability appears when $\mu_{\rm eff}^{2}$ is negative, and:
$\frac{d^{2}f(\phi)}{d\phi^{2}}\Bigg{|}_{\phi=0}\neq 0\quad\text{or}\quad\frac{d^{2}g(\phi)}{d\phi^{2}}\Bigg{|}_{\phi=0}\neq 0.$ (25)
When $\mu_{\rm eff}^{2}<0$ a new branch of scalarized solutions bifurcate from the original scalar-free solutions (GR black holes). These scalarized solutions form a family of solutions which are continuously connected to the scalar-free solutions and characterized by additional parameters (e.g., the scalar field’s value at the horizon). The quantities $\mathcal{I}_{GB}$ and $\mathcal{I}_{M}$ influence both the onset of stability and the resulting scalarized solutions.
The choice of the coupling functions $f(\phi)$ and $g(\phi)$ plays an important role. Various forms have been proposed in the literature, which satisfy the condition that $f(0)=0$, $g(0)=1$, so that for vanishing scalar field one recovers the Einstein-Maxwell action, whose unique spherically symmetric solution is the Reissner-Nordström (RN) metric:
Functions for $f(\phi)$ [44, 84]:
  1. 1.
    $\frac{\eta\phi^{2}}{2}$
  2. 2.
    $\frac{\eta}{2\beta}(1-e^{-\beta\,\phi^{2}})$
  3. 3.
    $\frac{\eta}{\beta^{2}}(1-\frac{1}{\cosh(\beta\,\phi)})$
  4. 4.
    $\frac{\eta\phi^{2}}{2\,(1+\beta^{2}\,\phi^{2})}$
Functions for $g(\phi)$, [45, 69, 72]:
  1. 1.
    $1-\alpha\,\phi^{2}$
  2. 2.
    $e^{-\alpha\,\phi^{2}}$
  3. 3.
    $\cosh(\sqrt{2|\alpha|}\,\phi)$
  4. 4.
    $\frac{1}{1+\alpha\phi^{2}}$
For small $\phi$ they all have the same behavior, so for simplicity here we adopt the simplest quadratic form :
$f(\phi)=\frac{\eta}{2}\phi^{2},\qquad g(\phi)=1-\alpha\phi^{2}$ (26)
[…]
If we take the ansatz (9) and put
$N(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\ \ ,\qquad\sigma(r)=1,\qquad V(r)=\frac{Q}{r}\ .$ (27)
we get the metric of the RN black hole, which has an event horizon located at
$r_{H}=M+\sqrt{M^{2}-Q^{2}}.$ (28)
For this RN background, the matter and the Gauss-Bonnet invariants are
$\mathcal{I}_{M}=-\frac{2Q^{2}}{r^{4}}<0\ ,\quad\text{and}$ (29)
$\mathcal{I}_{GB}=\frac{8}{r^{8}}\left[6M^{2}r^{2}-12Q^{2}Mr+5Q^{4}\right]$ (30)
The total effective mass becomes
$\mu_{\rm eff}^{2}=\frac{-2\eta}{r^{8}}\left(6M^{2}r^{2}-12Q^{2}Mr+5Q^{4}\right)\ +\frac{\alpha Q^{2}}{r^{4}}$ (31)
We will analyze the stability of the RN solution within this theory (we neglect back-reaction). This approach provides scalarization thresholds and identifies the onset of tachyonic instability associated with spontaneous scalarization.
Let us write the linearized KG equation (scalar field perturbation equation) on a fixed RN background, then decompose the scalar field perturbations into real spherical harmonics
$\delta\phi=U(r)Y_{\ell m}(\theta,\varphi)$ (32)
This allows the scalar field perturbations equation (linearized KG equation) to be only radially dependent and written in the form
$\left[r^{2}N(r)U^{\prime}(r)\right]^{\prime}=U(r)(\ell(\ell+1)-\frac{\eta}{r^{2}}\left(\frac{12M^{2}}{r^{2}}+\frac{10Q^{4}}{r^{4}}-\frac{24MQ^{2}}{r^{3}}\right)+\alpha\frac{Q^{2}}{r^{2}}),$ (33)
In the analysis of the KG equation in the theory (1), the spherical symmetry make it easy to separate the wave equation (characterized by $\ell=0$).
For $l=0$ this equation can be written as
$(Q^{2}+r(-2M+r))(U^{\prime\prime})(r)=\frac{\left(-10Q^{4}\eta-12M^{2}r^{2}\eta+Q^{2}(r^{4}\alpha+24Mr\eta)\right)U(r)}{r^{6}}+2(M-r)U^{\prime}(r).$ (34)
To facilitate our investigation, we introduce dimensionless variables :
$q\equiv\frac{Q}{M},\quad u\equiv\frac{U}{M},\quad\rho\equiv\frac{r}{M},\quad\gamma\equiv\frac{\eta}{M^{2}}$ (35)
leading to the dimensionless form of the equation:
$\rho^{6}((q^{2}+(-2+\rho)\rho)u^{\prime\prime}(\rho)+2(-1+\rho)u^{\prime}(\rho))=(q^{2}\rho^{4}\alpha-2(5q^{4}-12q^{2}\rho+6\rho^{2})\gamma)u(\rho).$ (36)
The differential equation (36) can be solved numerically. We start our numerical integration near the event horizon (at $\rho=\rho_{H}+10^{-5}$), with $\rho_{H}=1+\sqrt{1-q^{2}}$  [^5] We will integrate the equation for the range of values $0\leqslant q\leqslant 1$; $q=1$ is the extremal RN black holes. using a power series solution for $u$ valid in that region. The leading-order behavior near the horizon is given by:
$u=u_{H}+\frac{(q^{2}\rho_{H}^{4}\alpha+\gamma(-10q^{4}-12\rho_{H}^{2}+q^{2}24\rho_{H}))u_{H}}{2(\rho_{H}-1)\rho_{H}^{6}}(\rho-\rho_{H})$ (37)
We extend the integration to a large value of $\rho$, typically around $10^{5}$, for specific values of the coupling constants $(\gamma,\,\alpha)$ and an arbitrary horizon value $u_{H}$. For every set of these parameters, the equation (36) yields a distinct solution. Our focus is on solutions $u$ that vanish at spatial infinity.
Our exploration covers a parameter space defined by the reduced charge $q$ and the coupling parameters $(\gamma,\,\alpha)$. This results in a three-dimensional domain characterizing the model in question. To simplify our analysis, we fix one of the coupling constants, for instance $\alpha$, and vary the other, $\gamma$.
We varied these parameters systematically to map out the region in the parameter space where the GR is unstable in order to identify where scalarization can happen. This region is characterized by the onset of tachyonic instability, indicated by the conditions on the effective mass squared, $\mu_{\rm eff}^{2}$, as outlined in Eq. (23).
Figure 1: This figure illustrates the behavior of the lower limit of the dimensionless Gauss-Bonnet coupling constant $\gamma$, which is necessary to guarantee the instability of the Reissner-Nordström black hole. It showcases the dimensionless ratio $\gamma=\eta/M^{2}$ for a set of Reissner-Nordström solutions that support the $\ell=0$ scalar cloud, plotted against the reduced electric charge $q=Q/M$.The inset is a magnification of the main plot near the region where $\eta$ becomes negative
In Fig. 1 we plot, for six fixed values of the matter coupling $\alpha$, the minimum dimensionless Gauss-Bonnet strength $\gamma=\eta/M^{2}$ that triggers the instability of a Reissner–Nordström black hole as a function of the charge ratio $q=Q/M$. The six panels are arranged by $\alpha$ from the top left $(\alpha=0)$ to the bottom right $(\alpha=-40)$. Each curve therefore marks the bifurcation line where a non-trivial scalar configuration first appears. The bullet list that follows comments on the behavior in every panel.
  • Top-left, $\alpha=0$. The curve reproduces the results already discussed in [84, 86].
  • Top-right, $\alpha=-2$ and $\alpha=-5$. The blue $(\alpha=-2)$ and orange $(\alpha=-5)$ curves bend below the $\gamma=0$ axis at high charge. The Maxwell coupling therefore allows scalarization with $\eta<0$ once $q$ is large.
  • Middle-left, $\alpha=-10$. The fundamental branch crosses $\gamma=0$ at $q\approx 0.6$, extending the $GB^{-}$ window that earlier work found only near $q\approx 0.96$ [85, 86]. Other branches appear, numerical checks confirm that the curve that appears just above the first fundamental branch is the first excited scalar mode with one radial node $(n=1)$.
[…]
  • Bottom-left, $\alpha=-30$. The fundamental branch turns negative already around $q\approx 0.3$. Three excited branches track higher-node clouds.
  • Bottom-right, $\alpha=-40$. Almost the entire plot lies below $\gamma=0$. Scalarization with negative $\eta$ is now possible for moderate charges $(q\gtrsim 0.2)$ and several higher-node branches follow the same trend.
Overall, increasing $|\alpha|$ widens the parameter space in which the instability occurs and introduces additional node branches. It permits $\eta<0$ solutions with less restrictive conditions. The critical value $q^{(-)}$ at which $\eta$ flips sign decreases monotonically with $|\alpha|$; larger couplings push $\eta<0$ scalarization window down to $q\approx 0.2$, making negative-$\eta$ scalarization accessible at a much broader class of RN BH.
Earlier studies located this so-called $GB^{-}$ branch in narrow regimes. In case of spherical symmetry studied in [85, 86] RN clouds with $\eta<0$ were found only near extremality $q\approx 0.957058$. For rotating black holes, [148, 149, 150] reported the so-called spin-induced scalarization with $\eta<0$ scalarization when the spin is high ($J/M^{2}\gtrsim 0.5$. In both cases the Gauss-Bonnet invariant turns negative.
III Numerical scalarized black hole solutions and their properties
III.1 Numerical procedure
We solve the system of ordinary differential equations in (11) with the boundary conditions in (12) and (14). These equations describe a boundary value problem with conditions at the event horizon and at large distance. Standard scaling arguments allow to set the horizon radius to $r_{H}=1$.
We work with three parameters: $\eta$, $\alpha$, and $Q$. We fix $\alpha$ and load from the bifurcation data an initial list of values of ($q,\eta/M^{2}$) that satisfy a certain ratio $Q/\sqrt{\eta}$, which defines the onset of the instability of the RN background. Starting from the bifurcation (GR) value, we vary $\eta$ in small increments or decrements while keeping $Q/\sqrt{\eta}$ constant. For each step, we solve for the scalar field value at the horizon so that the field vanishes at large distance. This strategy leads to families of asymptotically flat black holes with scalar hair.
We adopt a shooting method combined with a stiff solver. We start the integration near $r=r_{H}+10^{-4}$, using the near-horizon expansions in (12). We move outward to $r\sim 10^{3}$ and compare the result to the far-field expansions in (14). We adjust the shooting parameters, which include $\phi_{0}$ and $\sigma_{0}$, until the scalar field approaches zero at large $r$ and the metric converges to a RN form. We then rescale $\sigma$ at large $r$ to fix the normalization. In practice, we set $\sigma_{0}=1$ at the start, then adjust it to $\sigma_{0}=1/\sigma_{\infty}$.
When the step in $\eta$ fails to produce a well-behaved solution, we revert to the previous successful solution and reduce the step size. We require $\Delta>0$ and the solutions for mass and charge to remain real. The local integration error is set close to $10^{-10}$.
This procedure generates black hole solutions labeled by $\alpha$, $\eta$, and $Q/\sqrt{\eta}$. They arise from the Reissner–Nordström configuration through a tachyonic instability.
III.2 Scalarized black hole solutions
In the following, we focus only on the case where $\eta$ is positive and $\alpha$ negative (which is the usual choice of sign for scalarization both in EsGB and EMS). One of the most important quantities to study for scalarized black holes is the scalar charge, which gives a clear idea of how scalarized black holes bifurcate from the GR ones and how they evolve.
Figure 2: Normalized scalar charge $D/\sqrt{\eta}$ with respect to the normalized mass $M/\sqrt{\eta}$ for a fixed normalized electric charge $Q/\sqrt{\eta}$, comparing different values of $\alpha$, for fixed values of $Q/\sqrt{\eta}$
Figure 2 shows the plots of the scalar charge $\frac{D}{\sqrt{\eta}}$ with respect to the mass $\frac{M}{\sqrt{\eta}}$ for fixed ratio $Q/\sqrt{\eta}$, comparing several solutions with different $\alpha$. The scalar charge $\frac{D}{\sqrt{\eta}}$ versus mass $\frac{M}{\sqrt{\eta}}$ plots for scalarized black holes show that each choice of the coupling parameter $\alpha$ yields a distinct branch of solutions bifurcating from RN at different threshold mass. In single-coupling models (pure EsGB or EsGB plus a Ricci term) those mass thresholds are identical. In the present model, the threshold mass depends strongly on $\alpha$ and on $Q/\sqrt{\eta}$ This allows a shift of the bifurcation mass $\tilde{M}_{th}$ over a wide interval: the largest onset masses are up to three to four times the smallest ones. Furthermore after the bifurcation point, some scalarized branches extend toward higher masses while others extend toward lower masses. The direction of this extension appears to correlate with stability.
  • Varying threshold masses: Scalarization in this model sets in at different threshold masses $\tilde{M}_{th}=\frac{M}{\sqrt{\eta}}$ for different $\alpha$ values. This is in contrast with simpler cases of pure EsGB or EsGB with Ricci coupling, where all the branches share the same onset mass. For $\alpha=0$ (no Maxwell coupling), the threshold masses $\tilde{M}_{th}$ are very close and almost identical (small difference), as previously found in [84]. By introducing a Maxwell coupling, the scalarization onset spans a much wider range of masses (some black holes have $\tilde{M}_{th}$ four times the lightest one), with each $\alpha$ producing its own distinct threshold.
  • Branch behavior at low charge: At low charge fraction $Q/\sqrt{\eta}=0.2$, the direction in which scalarized branches extend after bifurcation depends on $\alpha$ Coupling values $\alpha=0,-2,-5$ produce branches that extend toward higher mass values beyond the threshold (rightward on the plot). In contrast $\alpha=-20,-30,-40$ yield branches extending toward lower masses. The case $\alpha=-10$ shows a critical turning point where the branch has a curve that changes direction when analyzed individually in a small interval of $M/\sqrt{\eta}$ (see Fig. 3).
    Figure 3: Scalar-charge branch for $\alpha=-10$ and $Q/\sqrt{\eta}=0.2$.
  • Branch behavior at higher charge: For larger charge to coupling ratio (e.g. $Q/\sqrt{\eta}=1.6$ or above), all the scalarized solutions branches tend to extend toward smaller masses.
  • Stability implications: The direction in which a scalarized branch evolves is believed to indicate its stability. Branches that extended toward smaller masses (decreasing $M/\sqrt{\eta}$ as the scalar charge grows) are likely stable solutions, whereas those that extend toward larger masses beyond the bifurcation point may be unstable. This interpretation is in line with suggestions of [90, 91], which means that introducing the Matter coupling may have a stabilizing effect.
  • Convergent endpoints: Several of the scalar charge curves for different $\alpha$ converge toward similar end points in the plane. Despite their different onset masses and trajectories, many branches approach nearly the same final scalar charge (and associated mass) at their termination. This suggests that different coupling strengths $\alpha$ can lead to scalarized black holes with comparable limiting properties.
III.3 Thermodynamic properties
We analyze the thermodynamic properties of the scalarized black hole solutions. Key quantities include the horizon area and the Hawking temperature. For the metric ansatz (9), the area and the temperature are :
$A_{H}=4\pi r_{H}^{2},\qquad T_{H}=\frac{\sigma(r_{H})N^{\prime}(r_{H})}{4\pi}.$ (38)
The black hole entropy $S_{H}$ in this model presents corrections to the Bekenstein-Hawking law, due to the non-minimal coupling with the Gauss-Bonnet term. We adopt the Wald–Iyer approach [151, 152], which associates a Noether charge to the black hole entropy in theories with diffeomorphism-invariant Lagrangian: then the entropy $S_{H}$ is defined as
$S_{H}=2\pi\int_{H}\frac{\partial\mathcal{L}}{\partial R_{\mu\nu\alpha\beta}}\epsilon_{\mu\nu}\epsilon_{\alpha\beta},$ (39)
where $\mathcal{L}$ is the Lagrangian density and $\epsilon_{\alpha\beta}$ is the horizon’s volume form.
For the action in Eq. (1), this reduces to:
$S_{H}=\frac{1}{4}A_{H}+4\pi f(\phi_{0}),$ (40)
where $\phi_{0}$ is the scalar field value at the horizon. This expression holds for both static and rotating black holes .
For instance scalarized black holes in EMsGB theories satisfy Smarr relation [152, 153, 86]. For the static solutions considered here ($J=0,\Omega_{H}=0$), this relation is:
$M=2T_{H}S_{H}+\psi_{e}Q+M_{s}\ ,$ (41)
where $M_{s}$ is a contribution of the scalar field [154]:
$M_{s}=\frac{1}{4\pi}\int d^{3}x\sqrt{-g}\frac{f(\phi)}{f_{,\phi}(\phi)}\Box\phi.$
The first law of black hole thermodynamics in many of these models retains the form
$dM=T_{H}\,dS_{H}+\psi_{e}\,dQ,$ (42)
We can see that there is no direct contribution to this equation from the scalar charge $D$.
The most useful quantities to plot for studying black hole thermodynamic properties are the horizon area, the Hawking temperature, and the entropy. The entropy can be plotted as $S_{H}/\eta$ or as $S_{H}/S_{GR}$. The temperature and area can be represented either normalizing them with respect to the coupling constant, for instance $A_{H}/\eta$ and $T_{H}/\eta$, or by using the reduced quantities :
$a_{H}\equiv\frac{A_{H}}{16\pi M^{2}}\ ,\qquad t_{H}\equiv 8\pi T_{H}M=2MN^{\prime}(r_{H})\sigma(r_{H})\ ,$ (43)
Figure 4: The figures present several plots of the reduced area $a_{H}=\frac{A_{H}}{16\pi M^{2}}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$; each of them shows many solutions with different values of $\alpha$.
In Fig. 4 is reported the reduced area with respect to the reduced charge for several values of the other parameters: we can notice that the first two plots, corresponding to $Q/\sqrt{\eta}=0.2$ and 0.5473 show that the area of all the scalarized solutions is smaller than the RN one, while in the third plot, with $Q/\sqrt{\eta}=0.8$, the behavior of the solutions with $\alpha=-20,-30,-40$ is opposite, showing scalarized black holes with higher area than the RN one, which is in agreement with the fact that matter-induced scalarization becomes dominant due to the high value of the coupling to Maxwell ($\alpha$). From the value $Q/\sqrt{\eta}=1.6$ we clearly see that all the scalarized solutions with different $\alpha$ show higher area with respect to the RN one: this is expected even for the smallest value of $\alpha$ because of the high value of $Q/\sqrt{\eta}$ which enhances the contribution of scalarization induced by Maxwell in comparison with the one induced by the GB invariant.
Figure 5: The figures present several plots of the reduced temperature $t_{H}=8\pi MT_{H}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$; each one shows many solutions with different values of $\alpha$.
In Fig. 5 the reduced temperature of the scalarized solutions is plotted against the reduced charge. One can see that all scalarized solutions bifurcate from the GR branch, and have higher temperature than the RN black hole. Another important property is that they never reach extremality ($t_{H}$ never vanishes). Finally, it turns out that, starting from $Q/\sqrt{\eta}=1.6$, the scalarized black hole can become overcharged ($q>1$): this fact was never observed for $GB^{+}$ ($\eta>0$) scalarization before, but only in the case of $GB^{-}$ ($\eta<0$) as pointed out in [86].
Figure 6: The figures show the entropy of scalarized black holes normalized by the entropy of GR black holes with the same parameters ($S_{H}/S_{GR}$), with respect to the normalized mass $M/\sqrt{\eta}$: each one shows many solutions with different values of $\alpha$.
In Fig. 6 is depicted the entropy of scalarized black holes with respect to their GR counterpart. We see clearly that the entropy of all scalarized black holes is larger than the entropy of RN black holes, so they are thermodynamically preferred. For large values of $Q/\sqrt{\eta}$, the entropy is almost the same as in GR, especially for larger masses $M/\sqrt{\eta}$, while for small values of $Q/\sqrt{\eta}$ the entropy of scalarized black holes deviates from its GR value.
IV Conclusions
In this work, we introduced a new model of spontaneous black hole scalarization that includes couplings to both the Gauss-Bonnet curvature term and a $U(1)$ gauge field that can be taken as a Maxwell field type or a dark photon. By combining these curvature and matter couplings, we obtained charged black hole solutions with scalar hair that are not present in models with a single coupling. Our analysis revealed that the matter coupling extends the regime of scalarization even into regions with negative Gauss-Bonnet coupling, which was not possible with only curvature coupling. In practice, this means the scalar field can arise around a black hole under less restrictive conditions, thanks to the additional influence of the matter coupling.
The bifurcation plots in Fig. 1 show that the matter coupling allows scalarization with negative $\eta$ to occur with sub-extremal charges, which relaxes the near-extremality requirement observed in pure Gauss-Bonnet models associated with a Maxwell term.
Our numerical results further quantify how the black hole properties depend on the coupling constants. We examined the behavior of the scalar charge $D/\sqrt{\eta}$ as a function of the black hole mass $M/\sqrt{\eta}$. Generally, the scalar charge grows from zero at the threshold and then increases with $M$ for the scalarized solutions. The solutions show several key features. The presence of two coupling parameters leads to different mass thresholds for the onset of scalarization. We found that the critical mass at which the scalar field appears can vary widely: for certain values of the Gauss-Bonnet coupling $\eta$ and matter coupling $\alpha$, scalarization starts at much higher black hole masses than in the single-coupling cases, while for others it occurs earlier.
For different choices of $\alpha$ and $\eta$, the curves can follow different trajectories, but interestingly many branches tend towards similar values of scalar charge as they approach their endpoints.
By comparing the branches, we found indications of which Maxwell coupling may stabilize solutions with quadratic coupling as the Ricci coupling and self interactions did.
Finally, we found that our model permits overcharged black hole solutions (with charge $Q$ exceeding the extremal value $M$). We observed that beyond a certain charge (for example, $Q/\sqrt{\eta}\gtrsim 1.6$ in our examples), the scalarized branches can exceed the $q=1$ limit, yielding black holes with $q>1$. Such overcharged scalarized black holes were not seen in earlier Gauss-Bonnet-only models (except in the special case of negative $\eta$) and thus represent a novel feature introduced by the inclusion of the matter coupling.
Many future directions to extend this work are possible. One first step should derive the black hole solutions for $\eta<0$ ($GB^{-}$). The next one will be to study the stability of these solutions (radial and angular), the work of [102] make it simple to do the analysis. Future work can examine different coupling functions in our model. For instance, using an exponential form of coupling for one of the invariants or both of them instead of the quadratic forms.
One can also compute the quasi-normal modes of these black holes in order to test them with future detectors (LISA and Einstein Telescope), and study their shadow in light of the observations by the Event Horizon Telescope.
An ambitious direction should also target dynamical studies through numerical simulations for this model and see the impact of adding this new interaction to the hyperbolicity loss.
Another path that we suggest, is examining the implications of our findings on dark matter in the context of dark photon models. In a dark photon scenario, the Maxwell field in our model may represent a hidden $U(1)$ charge (a “dark” charge) carried by the black hole. Studying scalarization in this context may link our theoretical results to dark matter or dark sector phenomenology, and it could offer potential avenues to detect or constrain these exotic forms of black hole hair through their subtle gravitational or astrophysical signatures. A possible way is studying dark photon superradiance in this model [130, 131, 133, 135, 137, 136, 138].
Finally, it will be very interesting to verify if these theories can have any impact for explaining late time acceleration without breaking GW170817 constraints on the speed of gravitational waves [155, 156, 157, 158, 159].
Our new model with both curvature and matter couplings broadens the domain of spontaneous black hole scalarization and reveals a much richer structure of solutions. We have shown that scalarized black holes in this theory can exist under more flexible conditions, they appear to be more stable, possess higher entropy, and even achieve charge-to-mass ratios beyond the traditional extremal limit. These findings deepen our understanding of scalarization phenomena by introducing a new type of scalarization induced by two different sources and set the stage for further investigations into the stability and astrophysical implications of scalarized black holes.
Acknowledgments
We are very grateful to Pedro G.S. Fernandes, Hector O. Silva and Caio F.B. Macedo for discussions and comments.
Appendix A Field Equations of motion for the model
The expressions for $W$ and $Z$ are as follows:
$W$ $=16r(1-3N)^{2}f_{,\phi}(\phi)^{3}\phi^{\prime 2}\left(g(\phi)^{2}N(5N-1)\phi^{\prime 2}r^{4}+Q^{2}Ng_{,\phi}(\phi)\phi^{\prime}r-Q^{2}g(\phi)(N-1)\right)$
$\quad-4f_{,\phi}(\phi)^{2}\phi^{\prime}\left(-2Q^{4}+r^{3}(3N-1)(7N-1)g_{,\phi}(\phi)\phi^{\prime}Q^{2}\right)$
$\quad-4f_{,\phi}(\phi)^{2}\phi^{\prime}\left(2r^{2}g(\phi)\left(N\left(2r^{2}\phi^{\prime 2}-6r^{2}N\phi^{\prime 2}+2(N-1)(3N+1)f_{,\phi\phi}(\phi)\phi^{\prime 2}-9N+10\right)-1\right)Q^{2}\right)$
$\quad-4f_{,\phi}(\phi)^{2}\phi^{\prime}\left(2r^{4}g(\phi)^{2}N\left(2r^{2}\phi^{\prime 2}+N\left(r^{4}\phi^{\prime 4}-20r^{2}\phi^{\prime 2}+42r^{2}N\phi^{\prime 2}+6(N-1)^{2}\left(r^{2}\phi^{\prime 2}+2\right)f_{,\phi\phi}(\phi)\phi^{\prime 2}-3N+6\right)-3\right)\right)$
$\quad-2rf_{,\phi}(\phi)\left(Q^{4}+2r^{3}(1-4N)g_{,\phi}(\phi)\phi^{\prime}Q^{2}+2r^{2}g(\phi)\left(-r^{2}\phi^{\prime 2}+N\left(\left(5r^{2}-4(N-1)f_{,\phi\phi}(\phi)\right)\phi^{\prime 2}+3\right)-3\right)Q^{2}\right)$
$\quad-2rf_{,\phi}(\phi)\left(r^{4}g(\phi)^{2}\left(2r^{2}\phi^{\prime 2}\right.\right.$
$\quad\quad\left.\left.+N\left(-4(N-1)\left(r^{2}(3N-1)\phi^{\prime 2}+3(N-1)\right)f_{,\phi\phi}(\phi)\phi^{\prime 2}-2\left(r^{2}\phi^{\prime 2}+3\right)-N\left(r^{4}\phi^{\prime 4}+24r^{2}\phi^{\prime 2}-3\right)\right)+3\right)\right)$
$\quad-r^{5}\left(g_{,\phi}(\phi)Q^{2}+2rg(\phi)\phi^{\prime}\left(r^{2}g(\phi)\left(2(N-1)Nf_{,\phi\phi}(\phi)\phi^{\prime 2}+N+1\right)-Q^{2}\right)\right),$
$Z$ $=-2r^{2}g(\phi)N\left(8Q^{2}(N-1)f_{,\phi}(\phi)^{2}\left((3N+1)f_{,\phi}(\phi)\phi^{\prime}-r\right)+r^{2}g(\phi)\left(r^{5}+2f_{,\phi}(\phi)\left((1-7N)\phi^{\prime}r^{4}\right.\right.\right.$
$\quad\left.\left.\left.+6f_{,\phi}(\phi)\left(2r^{2}N(3N-1)\phi^{\prime 2}-(N-1)^{2}\right)r+4Nf_{,\phi}(\phi)^{2}\phi^{\prime}\left(6(N-1)^{2}+r^{2}(3N(2-5N)+1)\phi^{\prime 2}\right)\right)\right)\right).$

### v2

@@
-A new model of spontaneous scalarization of black holes induced by curvature and matter
+New model of spontaneous scalarization of black holes induced by curvature and matter
 Abstract
@@
 They might probe signatures of alternative gravitational theories which were not observed before.
-Black holes (BH) in GR are characterized by only three quantities: mass, spin and electric charge.
+Black holes (BHs) in GR are characterized by only three quantities: mass, spin and electric charge.
 No-hair theorems reinforce this claim and exclude extra scalar charges for some models when gravity couples to matter in the minimal way [10, 11].
@@
 Such extensions can emerge from fundamental physics, such as some attempts to unify gravity with other interactions naturally.
-A broad range of modified gravity models has been developed (see these review papers [15, 16, 17, 18, 19]), and upcoming strong-field tests offer a way to distinguish these alternative theories from GR in regimes where deviations may become pronounced.
+A broad range of modified gravity models has been developed (see the review papers [15, 16, 17, 18, 19]), and upcoming strong-field tests offer a way to distinguish these alternative theories from GR in regimes where deviations may become pronounced.
 One of the principal frameworks for modifying gravity is the class of scalar-tensor theories that introduce a dynamical scalar field alongside the tensor field of GR.
@@
 No-hair theorems were developed for some classes of scalar-tensor theories.
-Hawking proved it for Brans-Dicke theories [10], then [16] extend the proof to $f(R)$ theories.
+Hawking proved it for Brans-Dicke theories [10], then [16] extended the proof to $f(R)$ theories.
 Hui and Nicolis established no-hair theorems for most interesting classes of Horndeski theories [26].
@@
 In other words they admit "hairy" black holes solutions that evade no-hair theorems.
-As proved in these theorems scalar hair is forbidden in many scalar-tensor models unless specific conditions are met, as when the scalar is coupled to additional invariants like Gauss-Bonnet (GB) term [27].
+As proved in these theorems scalar hair is forbidden in many scalar-tensor models unless specific conditions are met, such as when the scalar is coupled to additional invariants like the Gauss-Bonnet (GB) term [27].
 The possibility of scalar hair on black holes marks a significant departure from standard GR and motivates further empirical tests in the strong gravity regime.
@@
 Hairy black hole solutions in this model were first found perturbatively in [35] and then numerically in [36].
-After that [27] showed that shift symmetric coupling of the scalar field to GB invariant allow to get hairy solutions.
-Recently [37, 38] discovered a new violation of no-hair theorems in EsGB theories for a large class of coupling functions including a quadratic coupling .
+After that Ref [27] showed that shift symmetric coupling of the scalar field to GB invariant allow hairy solutions.
+Recently, Refs [37, 38] discovered a new violation of no-hair theorems in EsGB theories for a large class of coupling functions including a quadratic coupling.
 EsGB theories provide a promising arena to explore strong-gravity phenomena and to confront theoretical predictions with observational data.
@@
 The dilaton coupling thereby alters the black hole’s charge-mass relation; BH can attain charge-to-mass ratios exceeding the extremal limit of RN black holes (they are overcharged).
-Models with scalar-Maxwell couplings, often inspired by high energy theory, thus provide a rich framework for how additional fields might influence electromagnetic interactions in strong gravity, offering distinctive black holes signatures that can be probed observationally.
+Models with scalar-Maxwell couplings, often inspired by high energy theory, thus provide a rich framework for how additional fields might influence electromagnetic interactions in strong gravity, offering distinctive black hole signatures that can be probed observationally.
 One of the most recent and actively studied mechanisms that allow compact objects to develop scalar hair only under some extreme conditions is spontaneous scalarization (see [41] for a review).
@@
 Later it was shown [43, 44] that spontaneous scalarization can apply to vacuum spacetimes (black holes) if the scalar couples to curvature or other matter invariants.
-Two broad classes of models realize this phenomenon: one is the curvature-induced scalarization through Gauss-Bonnet coupling [43, 44], the other is matter-induced scalarization through couplings to field such electromagnetism [45, 46].
+Two broad classes of models realize this phenomenon: one is the curvature-induced scalarization through Gauss-Bonnet coupling [43, 44], and the other is matter-induced scalarization through couplings to the field such as electromagnetism [45, 46].
 In both cases, a nonminimal coupling of a scalar field $\phi$ to an additional invariant (geometric invariant such as the GB term, the Ricci scalar or the Chern-Simons term [47, 48, 49], or matter invariant such as the electromagnetic tensor $F_{\mu\nu}F^{\mu\nu}$) triggers a tachyonic instability when the invariant is sufficiently large, causing the GR solution (with $\phi=0$) to become unstable, inducing a new ground state with $\phi\neq 0$.
@@
 $S_{\rm EsGB}=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\Big{[}R-2\nabla_{\mu}\phi\nabla^{\mu}\phi+f(\phi){\cal G}\Big{]}$~
-with $f(\phi)$ a coupling function.
+with $f(\phi)$ being a coupling function.
 The Schwarzschild black hole is a solution with $\phi=0$; however, if $f(\phi)$ is chosen such that $f^{\prime\prime}(0)\,{\cal G}>0$ induces a tachyonic mass for $\phi$, then above some coupling strength or below a certain mass scale the Schwarzschild solution becomes unstable.
@@
 This matter-induced scalarization is analogous to the Gauss–Bonnet case: the scalar hair develops only when the electric field (or charge-to-mass ratio $Q/M$) is sufficiently large, ensuring that minimally coupled GR solutions remain stable for weakly charged black holes.
-Such EMS models have been extensively studied [71, 72, 73, 73, 74, 75, 61, 76, 77, 78, 79, 80, 81, 82, 83] as a way to generate hairy black holes that are smooth deformations of the Reissner–Nordström solution, again obeying the principle that hair appears only in regimes where Maxwell field is above some threshold.
-There have been some attempts to study charged black holes with Gauss-Bonnet-induced scalarization, without any direct scalar-Maxwell coupling [84, 85, 86].
+Such EMS models have been extensively studied [71, 72, 73, 73, 74, 75, 61, 76, 77, 78, 79, 80, 81, 82, 83] as a way to generate hairy black holes that are smooth deformations of the Reissner–Nordström solution, again obeying the principle that hair appears only in regimes where the Maxwell field is above some threshold.
+There have been some studies of charged black holes with Gauss-Bonnet-induced scalarization, without any direct scalar-Maxwell coupling [84, 85, 86].
 These studies found that the inclusion of charge can enrich the scalarization behavior.
@@
 In particular, for the simple choice of a quadratic coupling function $f(\phi)\propto\phi^{2}$, all the resulting scalarized black hole solutions are unstable under radial (spherically symmetric) perturbations [88, 89].
-This means that although the hair can form, it does not represent a stable end state for evolution, small perturbations will drive the solution away (often back to the GR state).
+This means that although the hair can form, it does not represent a stable end state for evolution, and small perturbations will drive the solution away (often back to the GR state).
 More sophisticated coupling functions (for example, exponential [43, 88] or higher-order forms [90, 89]) can produce branches of solutions that are radially stable.
@@
 In fact, modes with higher angular momentum can become unstable on hairy solutions that are stable to monopole perturbations.
-For instance, in theories of EsGB with Ricci coupled quadratically to $\phi$ refereed as (EsGBR) [91, 92], a Quadrupole ($l=2$) instability was found by [93], which can be extended to EsGB theories.
+For instance, in theories of EsGB with Ricci coupled quadratically to $\phi$ referred to as (EsGBR) [91, 92], a quadrupole ($l=2$) instability was found by [93], which can be extended to EsGB theories.
 Another instability was reported in EsGB and in EsGBR, called gradient instability, where the sound speed of angular propagation becomes imaginary for even perturbations in the case of large harmonic index $l$ [94].
@@
 Motivated by these developments, in this work we propose a new model of black hole spontaneous scalarization induced by both curvature and matter, aiming to capture both curvature-induced and matter-induced scalarization within a single framework.
-The model introduces a simultaneous coupling of the scalar field to the Gauss–Bonnet invariant and to $U(1)$ gauge field (a dark photon or an electromagnetic field), incorporating the two mechanisms within one framework.
-Our aim is to leverage the two sources of scalarization to overcome some of the difficulties faced by the pure EsGB case.
+The model introduces a simultaneous nonminimal coupling of the scalar field to the Gauss–Bonnet invariant and to $U(1)$ gauge field (a dark photon or an electromagnetic field), incorporating the two mechanisms within one framework.
+One should have in mind that a theory which contains a nonminimal coupling to both Gauss-Bonnet and Maxwell terms is very well motivated.
+Earlier works have consider it as a low energy limit of string actions [32, 28, 30, 35], and black hole solutions were already investigated in this effective string theory [35, 102, 103, 104, 105].
+Our aim is to leverage the two sources of scalarization to try to overcome some of the difficulties faced by the pure EsGB case like instabilities.
 By including the matter-coupling directly, the scalar field has an additional channel to become tachyonic, which may widen the parameter space for stable scalarized solutions or provide new branches of hair that are absent in single-coupling models.
-Intuitively, the curvature coupling ensures that strong-gravity environments can trigger the scalar, while the matter coupling adds an extra contribution that may cure instabilities and modify the properties of the resulting black holes.
+Intuitively, the curvature coupling ensures that strong-gravity environments can trigger the scalar, while the matter coupling can also trigger scalarization and adds an extra contribution that modify the properties of the resulting black holes.
 This unified model is a logical next step because it reduces to the known cases in the appropriate limits (recovering pure Gauss-Bonnet scalarization when charge is zero, and pure Maxwell-induced scalarization when the Gauss-Bonnet coupling is turned off).
@@
 Another motivation, comes from the fact that this model can be seen as a subset of Maxwell-Horndeski theories, whose linear stability has already been studied [106].
-This will allow us to check easily if the model suffers from the same instabilities reported previously.
-We should add that by analogy to similar cases this model might be seen as a low energy limit of string actions [35].
+This will allow us to easily check if the model suffers from the same instabilities reported previously.
 Finally, let us clarify the interpretation of the $U(1)$ gauge field.
@@
 In the following sections, we detail the construction of this model with two sources and explore its black hole solutions and their properties.
-Section II introduces the model and field equations, then analyzes the onset of scalarization, Section III presents numerical solutions of the scalarized black holes their physical properties like the scalar charge and thermodynamic properties, and Section IV discusses conclusions and possible future directions.
-In this study, we adopt the convention $4\pi G=1=4\pi\epsilon_{0}$ for ease of calculation.
-The spacetime signature is (,+,+,+)(-,+,+,+)( - , + , + , + ).
+Sec II introduces the model and field equations, then analyzes the onset of scalarization, Sec III presents numerical solutions of the scalarized black holes their physical properties like the scalar charge and thermodynamic properties, and Sec IV discusses conclusions and possible future directions.
+In this study, we adopt the convention $4\pi G=1=4\pi\epsilon_{0}$ for convenience.
+The spacetime signature is (,+,+,+)(-,+,+,+).
 Our analysis is limited to spherically symmetric solutions, meaning that all metric and matter fields are functions of the radial coordinate only.
@@
 II.1 Description of the model
-We consider a family of theories [^2] These theories can be considered as a generalization of Einstein-Maxwell-scalar-Gauss-Bonnet theories, and a subclass of Maxwell-Horndeski theories.
+We consider a family of theories [^2] These theories can be considered as a generalization of Einstein-Maxwell-scalar-Gauss-Bonnet theories (EMsGB), and a subclass of Maxwell-Horndeski theories.
 where a real scalar field couples non-minimally both to the Gauss-Bonnet invariant and to a $U(1)$ gauge field (e.g.
 an electromagnetic field).
-The action is :
+The action is
 $S=\frac{1}{16\pi}\int d^{4}x\,\sqrt{-g}\Big{[}R-2\nabla_{\mu}\phi\nabla^{\mu}\phi+f(\phi)\mathcal{G}-g(\phi)F_{\mu\nu}F^{\mu\nu}\Big{]}\ ,$ (1)
-with $R$ the Ricci scalar and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ where $A_{\mu}$ is a $U(1)$ gauge field, $F_{\mu\nu}$ is its strength tensor (that can be identified as a Maxwell tensor), $\phi$ a scalar field non-minimally coupled to gravity and $U(1)$ gauge field through the functions $f(\phi)$ and $g(\phi)$ respectively, and $\mathcal{G}$ is the Gauss-Bonnet invariant which is defined as follows:
+with $R$ the Ricci scalar and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ where $A_{\mu}$ is a $U(1)$ gauge field, $F_{\mu\nu}$ is its strength tensor (that can be identified as a Maxwell tensor), $\phi$ is a scalar field non-minimally coupled to gravity and the $U(1)$ gauge field through the functions $f(\phi)$ and $g(\phi)$ respectively, and $\mathcal{G}$ is the Gauss-Bonnet invariant which is defined as follows:
 $\mathcal{G}=R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}\ .$ (2)
@@
     and $\varepsilon_{\alpha\beta\rho\sigma}$ is the Levi-Civita tensor.
-  • Klein-Gordon equation:
-    $\Box\phi+\frac{1}{4}f_{,\phi}(\phi)\mathcal{G}-\frac{1}{4}g_{,\phi}(\phi)F_{\mu\nu}F^{\mu\nu}=0$ (7)
+  • Klein-Gordon (KG) equation:
+    $\Box\phi+\frac{1}{4}f_{,\phi}(\phi)\mathcal{G}-\frac{1}{4}g_{,\phi}(\phi)F_{\mu\nu}F^{\mu\nu}=0\$(7)
   • Maxwell’s equations:
@@
 $ds^{2}=-N(r)\sigma(r)^{2}dt^{2}+\frac{dr^{2}}{N(r)}+r^{2}\big{(}d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}\big{)}\ ,~~{\rm with}~~N(r)\equiv 1-\frac{2\,m(r)}{r},$ (9)
-where $m(r)$ is the Misner-Sharp mass function [147] and $\sigma(r)$ a second metric function.
-Spherical symmetry, in the absence of a magnetic charge [^3] A magnetic charge would also be compatible with spherical symmetry, but it shall not be considered here - see [70] for dyonic charged BHs in the context of EMS theories., imposes an electrostatic 4-vector potential, $A=V(r)\,dt$, and a scalar field that depends only on $r$, $\phi=\phi(r)$.
-Substituting into the field equations leads to a set of three coupled differential equations for the radial functions $N$, $\sigma$, and $\phi$, with an additional constraint.
-This system of equations can be put in a form suitable for numerical integration:
-$N^{\prime}=F_{1}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),\qquad$ $\sigma^{\prime}=F_{2}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),$ (10)
-$\phi^{\prime\prime}=F_{3}(N,\sigma,\phi,\phi^{\prime}),\quad$ $(\frac{V^{\prime}g(\phi)r^{2}}{\sigma})^{\prime}=0.$
+where $m(r)$ is the Misner-Sharp mass function [151] and $\sigma(r)$ is a second metric function.
+Spherical symmetry, in the absence of a magnetic charge [^3] A magnetic charge would also be compatible with spherical symmetry, but it shall not be considered here - see [70] for dyonic charged BHs in the context of EMS theories., imposes an electrostatic 4-vector potential, A=V(r)dtA=V(r)\,dt, and a scalar field that depends only on $r$, $\phi=\phi(r)$.
+Substituting this ansatz into the field equations leads to a set of three coupled differential equations for the radial functions $N$, $\sigma$, and $\phi$, with an additional constraint.
+This system of equations can be put in a form suitable for numerical integration as
+$N^{\prime}=F_{1}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),\qquad$ $\sigma^{\prime}=F_{2}(N,\sigma,\phi,\phi^{\prime},\phi^{\prime\prime}),$ $\phi^{\prime\prime}=F_{3}(N,\sigma,\phi,\phi^{\prime})$ (10)
+$\qquad\text{and}\quad(\frac{V^{\prime}g(\phi)r^{2}}{\sigma})^{\prime}=0,$
 In more detail, these equations can be written as
@@
 II.3 Boundary conditions and physical quantities of interest
-To solve the set of four coupled ordinary differential equations (11), one must implement the appropriate boundary conditions.
-We impose regularity at the horizon, $r=r_{H}$, the field equations can be approximated by a power series expansion in $r-r_{H}$ as
+To solve the set of four coupled ordinary differential equations Eq (11), one must implement the appropriate boundary conditions.
+We impose regularity at the horizon, $r=r_{H}$, and the field equations can be approximated by a power series expansion in $r-r_{H}$ as
 $N=N_{0}+N_{1}(r-r_{H})+\cdots\ \ ,\qquad\sigma=\sigma_{0}+\sigma_{1}(r-r_{H})+\cdots\ ,$ (12)
 $\phi=\phi_{0}+\phi_{1}(r-r_{H})+\cdots\ ,\qquad\ \ V=V_{1}(r-r_{H})+\cdots\ ,$
-where the coefficients are given by
+where the coefficients are obtained by using the expansion into the field equations:
 $N_{0}$ $=0,$ $N_{1}$ $=\frac{r_{H}^{2}g(\phi_{0})-Q^{2}}{r_{H}^{2}g(\phi_{0})\left(2\phi_{1}f_{,\phi}(\phi_{0})+r_{H}\right)},$ (13)
@@
 $A$ $=4r_{H}f_{,\phi}(\phi_{0})\left(Q^{2}r_{H}^{2}f_{,\phi}(\phi_{0})g_{,\phi}(\phi_{0})-Q^{2}g(\phi_{0})\left(4f_{,\phi}(\phi_{0})^{2}+r_{H}^{4}\right)+r_{H}^{6}g(\phi_{0})^{2}\right).$
-The coefficients depend on the two parameters $\phi_{0}$ and $\sigma_{0}$, where the subscript 00 denotes functions evaluated at the horizon $r_{H}$.
+The coefficients depend on the two parameters $\phi_{0}$ and $\sigma_{0}$, where the subscript 0 denotes functions evaluated at the horizon $r_{H}$.
 The equation determining $\phi_{1}$ is obtained from the scalar field equation near the horizon and has two solutions, but we choose the only one that gives a black hole solution $\phi_{1}=\frac{B+\sqrt{\Delta}}{A}$, while the other solution $\phi_{1}=\frac{B-\sqrt{\Delta}}{A}$ does not lead to an asymptotically flat BH [36].
 To obtain a physical solution we need to ensure regularity by imposing that $\Delta\geq 0$ which keeps $\phi_{1}$ real and $\phi^{\prime\prime}_{0}$ finite [36].
-At spatial infinity we expand in power series of $1/r$, so the leading order expansion for far fields is
+At spatial infinity we expand in power series of $1/r$, so the leading-order expansion for far fields is
 $N\approx 1-\frac{2M}{r},\qquad\sigma\approx 1-\frac{D^{2}}{2r^{2}},\qquad$ $\phi\approx\frac{D}{r},\qquad\ \ V\approx\psi_{e}+\frac{Q}{r}$ (14)
@@
 $+\frac{r^{4}\!\bigl{(}3N^{\prime}(r)\sigma^{\prime}(r)+\sigma(r)N^{\prime\prime}(r)+2N(r)\sigma^{\prime\prime}(r)\bigr{)}^{2}}{r^{4}\sigma(r)^{2}}.$
-Scaling symmetry
+II.4 Scaling symmetry
 The field equations of the model exhibit a scaling symmetry [86, 44, 56], so that the solutions are invariant under the rescaling transformations
 $r\to\beta r\,,\quad M\to\beta M\,,\quad\eta\to\beta^{2}\eta\,,\quad Q\;\to\;\beta\,Q,$ (17)
-where $\beta>0$ is an arbitrary scaling parameter, $\eta$ is the coupling constant of the GB term, defined so that $f=\eta\,\tilde{f}$, with dimensionless $\tilde{f}$, so that $\eta$ has dimension of length square.
+where $\beta>0$ is an arbitrary scaling parameter, and $\eta$ is the coupling constant of the GB term, defined so that $f=\eta\,\tilde{f}$, with dimensionless $\tilde{f}$, so that $\eta$ has the dimension of length square.
 This symmetry and freedom in choosing the length scale implies that physical quantities of the black hole solution should be expressed in terms of dimensionless combinations that remain invariant under these transformations.
-To characterize the black hole solutions we introduce the reduced quantities :
+To characterize the black hole solutions we introduce the reduced quantities
 $q=\frac{Q}{M}\ ,\quad\frac{\eta}{M^{2}},\quad\frac{D}{\sqrt{\eta}},\quad\frac{Q}{\sqrt{\eta}},\quad\frac{M}{\sqrt{\eta}},\quad\frac{\eta}{r_{H}^{2}}$ (18)
@@
 All dimensional quantities presented in the results are normalized accordingly, either by appropriate powers of $M$ or $\sqrt{\eta}$ or $r_{H}$.
-II.4 Conditions for scalarization: Scalarization threshold and coupling functions
+II.5 Conditions for scalarization: Scalarization threshold and coupling functions
 The scalar field $\phi$ obeys the Klein-Gordon equation
@@
 $\mathcal{I}_{M}=F_{\mu\nu}F^{\mu\nu}\ ,\quad\mathcal{I}_{GB}=\mathcal{G}.$ (20)
-Spontaneous scalarization occurs when the theory admits solutions with $\phi=0$. This demands
+Two conditions are required for spontaneous scalarization to happen, as follows:
+  • When the theory admits solutions with $\phi=0$. This demands
     $\frac{df}{d\phi}\Big{|}_{\phi=0}=0,\quad\frac{dg}{d\phi}\Big{|}_{\phi=0}=0.$ (21)
     These conditions are generally met in models with $\mathbb{Z}_{2}$ symmetry (under $\phi\rightarrow-\phi$) and allow the existence of standard electro-vacuum GR solutions.
-A transition from a scalar-free black hole to one that supports a nontrivial scalar can occur if the second derivatives of $f(\phi)$ and $g(\phi)$ near $\phi=0$ create a tachyonic instability.
-Small-$\phi$ expansions for both coupling functions show:
+  • If the second derivatives of $f(\phi)$ and $g(\phi)$ near $\phi=0$ create a tachyonic instability. Small-$\phi$ expansions for both coupling functions show
     $f(\phi)=f(0)+\frac{1}{2}\frac{d^{2}f}{d\phi^{2}}\Big{|}_{\phi=0}\phi^{2}+\cdots,\quad g(\phi)=g(0)+\frac{1}{2}\frac{d^{2}g}{d\phi^{2}}\Big{|}_{\phi=0}\phi^{2}+\cdots,$ (22)
-The linearized the Klein-Gordon equation then becomes
+    The linearized Klein-Gordon equation then becomes
     $(\Box-\mu_{\rm eff}^{2})\delta\phi=0,\quad\text{where}\quad\mu_{\rm eff}^{2}=\mu_{\rm eff,f}^{2}+\mu_{\rm eff,g}^{2}$ (23)
@@
     $\mu_{\rm eff,f}^{2}=\frac{-1}{4}\frac{d^{2}f}{d\phi^{2}}\Big{|}_{\phi=0}\mathcal{I}_{GB},\quad\mu_{\rm eff,g}^{2}=\frac{1}{4}\frac{d^{2}g}{d\phi^{2}}\Big{|}_{\phi=0}\mathcal{I}_{M}$ (24)
-and $\mu_{\rm eff,f}^{2}$ and $\mu_{\rm eff,g}^{2}$ are the effective mass squares.
-A tachyonic instability appears when $\mu_{\rm eff}^{2}$ is negative, and:
+    and $\delta\phi$ is the scalar field perturbation, while $\mu_{\rm eff,f}^{2}$ and $\mu_{\rm eff,g}^{2}$ are the effective mass squares.
+    A tachyonic instability appears when $\mu_{\rm eff}^{2}$ is negative, and
     $\frac{d^{2}f(\phi)}{d\phi^{2}}\Bigg{|}_{\phi=0}\neq 0\quad\text{or}\quad\frac{d^{2}g(\phi)}{d\phi^{2}}\Bigg{|}_{\phi=0}\neq 0.$ (25)
-When $\mu_{\rm eff}^{2}<0$ a new branch of scalarized solutions bifurcate from the original scalar-free solutions (GR black holes).
+When $\mu_{\rm eff}^{2}<0$ a new branch of scalarized solutions may bifurcate from the original scalar-free solutions (GR black holes).
 These scalarized solutions form a family of solutions which are continuously connected to the scalar-free solutions and characterized by additional parameters (e.g., the scalar field’s value at the horizon).
@@
 The choice of the coupling functions $f(\phi)$ and $g(\phi)$ plays an important role.
-Various forms have been proposed in the literature, which satisfy the condition that $f(0)=0$, $g(0)=1$, so that for vanishing scalar field one recovers the Einstein-Maxwell action, whose unique spherically symmetric solution is the Reissner-Nordström (RN) metric:
+Various forms have been proposed in the literature, which satisfy the condition that $f(0)=0$, $g(0)=1$, so that for the vanishing scalar field one recovers the Einstein-Maxwell action, whose unique spherically symmetric solution is the Reissner-Nordström (RN) metric.
 Functions for $f(\phi)$ [44, 84]:
@@
     $\frac{\eta\phi^{2}}{2\,(1+\beta^{2}\,\phi^{2})}$
-Functions for $g(\phi)$, [45, 69, 72]:
+Functions for $g(\phi)$ [45, 69, 72]:
   1. 1.
@@
 $\mathcal{I}_{M}=-\frac{2Q^{2}}{r^{4}}<0\ ,\quad\text{and}$ (29)
-$\mathcal{I}_{GB}=\frac{8}{r^{8}}\left[6M^{2}r^{2}-12Q^{2}Mr+5Q^{4}\right]$ (30)
+$\mathcal{I}_{GB}=\frac{8}{r^{8}}\left[6M^{2}r^{2}-12Q^{2}Mr+5Q^{4}\right]\$(30)
 The total effective mass becomes
@@
 Let us write the linearized KG equation (scalar field perturbation equation) on a fixed RN background, then decompose the scalar field perturbations into real spherical harmonics
-$\delta\phi=U(r)Y_{\ell m}(\theta,\varphi)$ (32)
-This allows the scalar field perturbations equation (linearized KG equation) to be only radially dependent and written in the form
-$\left[r^{2}N(r)U^{\prime}(r)\right]^{\prime}=U(r)(\ell(\ell+1)-\frac{\eta}{r^{2}}\left(\frac{12M^{2}}{r^{2}}+\frac{10Q^{4}}{r^{4}}-\frac{24MQ^{2}}{r^{3}}\right)+\alpha\frac{Q^{2}}{r^{2}}),$ (33)
+$\delta\phi=U(r)Y_{\ell m}(\theta,\varphi)\$(32)
+This allows the scalar field perturbation equation (linearized KG equation) to be only radially dependent and written in the form
+$\left[r^{2}N(r)U^{\prime}(r)\right]^{\prime}=U(r)\left(\ell(\ell+1)-\frac{\eta}{r^{2}}(\frac{12M^{2}}{r^{2}}+\frac{10Q^{4}}{r^{4}}-\frac{24MQ^{2}}{r^{3}})+\alpha\frac{Q^{2}}{r^{2}}\right).$ (33)
 In the analysis of the KG equation in the theory (1), the spherical symmetry make it easy to separate the wave equation (characterized by $\ell=0$).
-For $l=0$ this equation can be written as
+For $\ell=0$ this equation can be written as
 $(Q^{2}+r(-2M+r))(U^{\prime\prime})(r)=\frac{\left(-10Q^{4}\eta-12M^{2}r^{2}\eta+Q^{2}(r^{4}\alpha+24Mr\eta)\right)U(r)}{r^{6}}+2(M-r)U^{\prime}(r).$ (34)
-To facilitate our investigation, we introduce dimensionless variables :
+To facilitate our investigation, we introduce dimensionless variables
 $q\equiv\frac{Q}{M},\quad u\equiv\frac{U}{M},\quad\rho\equiv\frac{r}{M},\quad\gamma\equiv\frac{\eta}{M^{2}}$ (35)
-leading to the dimensionless form of the equation:
+leading to the dimensionless form of the equation
 $\rho^{6}((q^{2}+(-2+\rho)\rho)u^{\prime\prime}(\rho)+2(-1+\rho)u^{\prime}(\rho))=(q^{2}\rho^{4}\alpha-2(5q^{4}-12q^{2}\rho+6\rho^{2})\gamma)u(\rho).$ (36)
-The differential equation (36) can be solved numerically.
-We start our numerical integration near the event horizon (at $\rho=\rho_{H}+10^{-5}$), with $\rho_{H}=1+\sqrt{1-q^{2}}$  [^5] We will integrate the equation for the range of values $0\leqslant q\leqslant 1$; $q=1$ is the extremal RN black holes.
+The differential equation Eq (36) can be solved numerically.
+We start our numerical integration near the event horizon (at $\rho=\rho_{H}+10^{-5}$), with $\rho_{H}=1+\sqrt{1-q^{2}}$  [^5] We will integrate the equation for the range of values $0\leqslant q\leqslant 1$, where $q=1$ is the extremal RN black hole.
 using a power series solution for $u$ valid in that region.
@@
 We extend the integration to a large value of $\rho$, typically around $10^{5}$, for specific values of the coupling constants $(\gamma,\,\alpha)$ and an arbitrary horizon value $u_{H}$.
-For every set of these parameters, the equation (36) yields a distinct solution.
+For every set of these parameters, Eq (36) yields a distinct solution.
 Our focus is on solutions $u$ that vanish at spatial infinity.
@@
 We varied these parameters systematically to map out the region in the parameter space where the GR is unstable in order to identify where scalarization can happen.
-This region is characterized by the onset of tachyonic instability, indicated by the conditions on the effective mass squared, $\mu_{\rm eff}^{2}$, as outlined in Eq.
+This region is characterized by the onset of tachyonic instability, indicated by the contribution of effective mass squared, $\mu_{\rm eff}^{2}$, as outlined in Eq.
 (33).
@@
 The six panels are arranged by $\alpha$ from the top left $(\alpha=0)$ to the bottom right $(\alpha=-40)$.
-Each curve therefore marks the bifurcation line where a non-trivial scalar configuration first appears.
-The bullet list that follows comments on the behavior in every panel.
+Each curve, therefore, marks the bifurcation line where a non-trivial scalar configuration first appears.
+The bullet list that follows comments on the behavior in every panel:
   • Top-left, $\alpha=0$. The curve reproduces the results already discussed in [84, 86].
@@
 The blue $(\alpha=-2)$ and orange $(\alpha=-5)$ curves bend below the $\gamma=0$ axis at high charge.
-The Maxwell coupling therefore allows scalarization with $\eta<0$ once $q$ is large.
+The Maxwell coupling, therefore, allows scalarization with $\eta<0$ once $q$ is large.
   • Middle-left, $\alpha=-10$.
 The fundamental branch crosses $\gamma=0$ at $q\approx 0.6$, extending the $GB^{-}$ window that earlier work found only near $q\approx 0.96$ [85, 86].
-Other branches appear, numerical checks confirm that the curve that appears just above the first fundamental branch is the first excited scalar mode with one radial node $(n=1)$.
-[…]
+Other branches appear, and numerical checks confirm that the curve that appears just above the first fundamental branch is the first excited scalar mode with one radial node $(n=1)$.
+  • Middle-right, $\alpha=-20$.
+The sign flip now occurs near $q\approx 0.4$.
+Three excited branches (n=1, 2, 3)(n=1,\;2,\;3) are visible.
+Larger $|\alpha|$, therefore, lowers the charge threshold and creates more bifurcation lines.
   • Bottom-left, $\alpha=-30$. The fundamental branch turns negative already around $q\approx 0.3$. Three excited branches track higher-node clouds.
@@
 It permits $\eta<0$ solutions with less restrictive conditions.
-The critical value $q^{(-)}$ at which $\eta$ flips sign decreases monotonically with $|\alpha|$; larger couplings push $\eta<0$ scalarization window down to $q\approx 0.2$, making negative-$\eta$ scalarization accessible at a much broader class of RN BH.
+The critical value $q^{(-)}$ at which $\eta$ flips sign decreases monotonically with $|\alpha|$, and larger couplings push the $\eta<0$ scalarization window down to $q\approx 0.2$, making negative-$\eta$ scalarization accessible at a much broader class of RN BH.
 Earlier studies located this so-called $GB^{-}$ branch in narrow regimes.
 In case of spherical symmetry studied in [85, 86] RN clouds with $\eta<0$ were found only near extremality $q\approx 0.957058$.
-For rotating black holes, [148, 149, 150] reported the so-called spin-induced scalarization with $\eta<0$ scalarization when the spin is high ($J/M^{2}\gtrsim 0.5$.
+For rotating black holes, Refs [152, 153, 154] reported the so-called spin-induced scalarization with $\eta<0$ scalarization when the spin is high ($J/M^{2}\gtrsim 0.5$.
 In both cases the Gauss-Bonnet invariant turns negative.
@@
 These equations describe a boundary value problem with conditions at the event horizon and at large distance.
-Standard scaling arguments allow to set the horizon radius to $r_{H}=1$.
+Standard scaling arguments allow us to set the horizon radius to $r_{H}=1$.
 We work with three parameters: $\eta$, $\alpha$, and $Q$.
@@
 III.2 Scalarized black hole solutions
-In the following, we focus only on the case where $\eta$ is positive and $\alpha$ negative (which is the usual choice of sign for scalarization both in EsGB and EMS).
+In the following, we focus only on the case where $\eta$ is positive and $\alpha$ is negative (which is the usual choice of sign for scalarization both in EsGB and EMS).
 One of the most important quantities to study for scalarized black holes is the scalar charge, which gives a clear idea of how scalarized black holes bifurcate from the GR ones and how they evolve.
@@
 Figure 2 shows the plots of the scalar charge $\frac{D}{\sqrt{\eta}}$ with respect to the mass $\frac{M}{\sqrt{\eta}}$ for fixed ratio $Q/\sqrt{\eta}$, comparing several solutions with different $\alpha$.
-The scalar charge $\frac{D}{\sqrt{\eta}}$ versus mass $\frac{M}{\sqrt{\eta}}$ plots for scalarized black holes show that each choice of the coupling parameter $\alpha$ yields a distinct branch of solutions bifurcating from RN at different threshold mass.
+The scalar charge $\frac{D}{\sqrt{\eta}}$ versus mass $\frac{M}{\sqrt{\eta}}$ plots for scalarized black holes show that each choice of the coupling parameter $\alpha$ yields a distinct branch of solutions bifurcating from RN at different threshold masses.
 In single-coupling models (pure EsGB or EsGB plus a Ricci term) those mass thresholds are identical.
-In the present model, the threshold mass depends strongly on $\alpha$ and on $Q/\sqrt{\eta}$ This allows a shift of the bifurcation mass $\tilde{M}_{th}$ over a wide interval: the largest onset masses are up to three to four times the smallest ones.
+In the present model, the threshold mass depends strongly on $\alpha$ and on $Q/\sqrt{\eta}$ .
+This allows a shift of the bifurcation mass $\tilde{M}_{th}$ over a wide interval: the largest onset masses are up to three to four times the smallest ones.
 Furthermore after the bifurcation point, some scalarized branches extend toward higher masses while others extend toward lower masses.
-The direction of this extension appears to correlate with stability.
+The direction of this extension appears to correlate with stability, as outlined below:
   • Varying threshold masses: scalarization in this model sets in at different threshold masses $\tilde{M}_{th}=\frac{M}{\sqrt{\eta}}$ for different $\alpha$ values.
@@
 By introducing a Maxwell coupling, the scalarization onset spans a much wider range of masses (some black holes have $\tilde{M}_{th}$ four times the lightest one), with each $\alpha$ producing its own distinct threshold.
-  • Branch behavior at low charge: At low charge fraction $Q/\sqrt{\eta}=0.2$, the direction in which scalarized branches extend after bifurcation depends on $\alpha$ Coupling values $\alpha=0,-2,-5$ produce branches that extend toward higher mass values beyond the threshold (rightward on the plot).
+  • Branch behavior at low charge: at low charge fraction $Q/\sqrt{\eta}=0.2$, the direction in which scalarized branches extend after bifurcation depends on $\alpha$ .
+Coupling values $\alpha=0,-2,-5$ produce branches that extend toward higher mass values beyond the threshold (rightward on the plot).
 In contrast $\alpha=-20,-30,-40$ yield branches extending toward lower masses.
@@
     Figure 3: Scalar-charge branch for $\alpha=-10$ and $Q/\sqrt{\eta}=0.2$.
-  • Branch behavior at higher charge: For larger charge to coupling ratio (e.g. $Q/\sqrt{\eta}=1.6$ or above), all the scalarized solutions branches tend to extend toward smaller masses.
+  • Branch behavior at higher charge: for larger charge to coupling ratio (e.g. $Q/\sqrt{\eta}=1.6$ or above), all the scalarized solution branches tend to extend toward smaller masses.
   • Stability implications: the direction in which a scalarized branch evolves is believed to indicate its stability.
 Branches that extended toward smaller masses (decreasing $M/\sqrt{\eta}$ as the scalar charge grows) are likely stable solutions, whereas those that extend toward larger masses beyond the bifurcation point may be unstable.
-This interpretation is in line with suggestions of [90, 91], which means that introducing the Matter coupling may have a stabilizing effect.
-  • Convergent endpoints: Several of the scalar charge curves for different $\alpha$ converge toward similar end points in the plane.
+This interpretation is in line with the suggestions of [90, 91], which means that introducing the matter coupling may have a stabilizing effect.
+  • Convergent end points: several of the scalar charge curves for different values of $\alpha$ converge toward similar end points in the plane.
 Despite their different onset masses and trajectories, many branches approach nearly the same final scalar charge (and associated mass) at their termination.
 This suggests that different coupling strengths $\alpha$ can lead to scalarized black holes with comparable limiting properties.
+The behavior of the curves suggests maybe the existence of an attractor.
 III.3 Thermodynamic properties
@@
 The black hole entropy $S_{H}$ in this model presents corrections to the Bekenstein-Hawking law, due to the non-minimal coupling with the Gauss-Bonnet term.
-We adopt the Wald–Iyer approach [151, 152], which associates a Noether charge to the black hole entropy in theories with diffeomorphism-invariant Lagrangian: then the entropy $S_{H}$ is defined as
+We adopt the Wald–Iyer approach [155, 156], which associates a Noether charge to the black hole entropy in theories with diffeomorphism-invariant Lagrangian: Then, the entropy $S_{H}$ is defined as
 $S_{H}=2\pi\int_{H}\frac{\partial\mathcal{L}}{\partial R_{\mu\nu\alpha\beta}}\epsilon_{\mu\nu}\epsilon_{\alpha\beta},$ (39)
@@
 where $\phi_{0}$ is the scalar field value at the horizon. This expression holds for both static and rotating black holes .
-For instance scalarized black holes in EMsGB theories satisfy Smarr relation [152, 153, 86]. For the static solutions considered here ($J=0,\Omega_{H}=0$), this relation is:
+For instance scalarized black holes in EMsGB theories satisfy the Smarr relation [156, 157, 86]. For the static solutions considered here ($J=0,\Omega_{H}=0$), this relation is
 $M=2T_{H}S_{H}+\psi_{e}Q+M_{s}\ ,$ (41)
-where $M_{s}$ is a contribution of the scalar field [154]:
+where $M_{s}$ is a contribution of the scalar field [158],
 $M_{s}=\frac{1}{4\pi}\int d^{3}x\sqrt{-g}\frac{f(\phi)}{f_{,\phi}(\phi)}\Box\phi.$
 The first law of black hole thermodynamics in many of these models retains the form
-$dM=T_{H}\,dS_{H}+\psi_{e}\,dQ,$ (42)
+$dM=T_{H}\,dS_{H}+\psi_{e}\,dQ.$ (42)
 We can see that there is no direct contribution to this equation from the scalar charge $D$.
@@
 The entropy can be plotted as $S_{H}/\eta$ or as $S_{H}/S_{GR}$.
-The temperature and area can be represented either normalizing them with respect to the coupling constant, for instance $A_{H}/\eta$ and $T_{H}/\eta$, or by using the reduced quantities :
+The temperature and area can be represented either by normalizing them with respect to the coupling constant, for instance $A_{H}/\eta$ and $T_{H}/\eta$, or by using the reduced quantities
 $a_{H}\equiv\frac{A_{H}}{16\pi M^{2}}\ ,\qquad t_{H}\equiv 8\pi T_{H}M=2MN^{\prime}(r_{H})\sigma(r_{H})\ ,$ (43)
-Figure 4: The figures present several plots of the reduced area $a_{H}=\frac{A_{H}}{16\pi M^{2}}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$; each of them shows many solutions with different values of $\alpha$.
+Figure 4: The figures present several plots of the reduced area $a_{H}=\frac{A_{H}}{16\pi M^{2}}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$, and each of them shows many solutions with different values of $\alpha$.
 In Fig.
-4 is reported the reduced area with respect to the reduced charge for several values of the other parameters: we can notice that the first two plots, corresponding to $Q/\sqrt{\eta}=0.2$ and 0.5473 show that the area of all the scalarized solutions is smaller than the RN one, while in the third plot, with $Q/\sqrt{\eta}=0.8$, the behavior of the solutions with $\alpha=-20,-30,-40$ is opposite, showing scalarized black holes with higher area than the RN one, which is in agreement with the fact that matter-induced scalarization becomes dominant due to the high value of the coupling to Maxwell ($\alpha$).
-From the value $Q/\sqrt{\eta}=1.6$ we clearly see that all the scalarized solutions with different $\alpha$ show higher area with respect to the RN one: this is expected even for the smallest value of $\alpha$ because of the high value of $Q/\sqrt{\eta}$ which enhances the contribution of scalarization induced by Maxwell in comparison with the one induced by the GB invariant.
-Figure 5: The figures present several plots of the reduced temperature $t_{H}=8\pi MT_{H}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$; each one shows many solutions with different values of $\alpha$.
+4 the reduced area is reported with respect to the reduced charge for several values of the other parameters.
+We can notice that the first two plots, corresponding to $Q/\sqrt{\eta}=0.2$ and 0.5473 show that the area of all the scalarized solutions is smaller than the RN one, while in the third plot, with $Q/\sqrt{\eta}=0.8$, the behavior of the solutions with $\alpha=-20,-30,-40$ is opposite, showing scalarized black holes with higher area than the RN one, which is in agreement with the fact that matter-induced scalarization becomes dominant due to the high value of the coupling to Maxwell ($\alpha$).
+From the value $Q/\sqrt{\eta}=1.6$ we clearly see that all the scalarized solutions with different values of $\alpha$ show higher area with respect to the RN one.
+This is expected even for the smallest value of $\alpha$ because of the high value of $Q/\sqrt{\eta}$ which enhances the contribution of scalarization induced by Maxwell in comparison with the one induced by the GB invariant.
+Figure 5: The figures present several plots of the reduced temperature $t_{H}=8\pi MT_{H}$ with respect to the reduced charge $q=Q/M$, for different fixed values of $Q/\sqrt{\eta}$, and each one shows many solutions with different values of $\alpha$.
 In Fig.
 5 the reduced temperature of the scalarized solutions is plotted against the reduced charge.
-One can see that all scalarized solutions bifurcate from the GR branch, and have higher temperature than the RN black hole.
+One can see that all scalarized solutions bifurcate from the GR branch, and have higher temperatures than the RN black hole.
 Another important property is that they never reach extremality ($t_{H}$ never vanishes).
-Finally, it turns out that, starting from $Q/\sqrt{\eta}=1.6$, the scalarized black hole can become overcharged ($q>1$): this fact was never observed for $GB^{+}$ ($\eta>0$) scalarization before, but only in the case of $GB^{-}$ ($\eta<0$) as pointed out in [86].
-Figure 6: The figures show the entropy of scalarized black holes normalized by the entropy of GR black holes with the same parameters ($S_{H}/S_{GR}$), with respect to the normalized mass $M/\sqrt{\eta}$: each one shows many solutions with different values of $\alpha$.
+Finally, it turns out that, starting from $Q/\sqrt{\eta}=1.6$, the scalarized black hole can become overcharged ($q>1$).
+This fact was never observed for $GB^{+}$ ($\eta>0$) scalarization before, but only in the case of $GB^{-}$ ($\eta<0$) as pointed out in [86].
+Figure 6: The figures show the entropy of scalarized black holes normalized by the entropy of GR black holes with the same parameters ($S_{H}/S_{GR}$), with respect to the normalized mass $M/\sqrt{\eta}$, and each one shows many solutions with different values of $\alpha$.
 In Fig.
-6 is depicted the entropy of scalarized black holes with respect to their GR counterpart.
+6, the entropy of scalarized black holes is depicted with respect to their GR counterparts.
 We see clearly that the entropy of all scalarized black holes is larger than the entropy of RN black holes, so they are thermodynamically preferred.
@@
 We found that the critical mass at which the scalar field appears can vary widely: for certain values of the Gauss-Bonnet coupling $\eta$ and matter coupling $\alpha$, scalarization starts at much higher black hole masses than in the single-coupling cases, while for others it occurs earlier.
-For different choices of $\alpha$ and $\eta$, the curves can follow different trajectories, but interestingly many branches tend towards similar values of scalar charge as they approach their endpoints.
-By comparing the branches, we found indications of which Maxwell coupling may stabilize solutions with quadratic coupling as the Ricci coupling and self interactions did.
+For different choices of $\alpha$ and $\eta$, the curves can follow different trajectories, but interestingly many branches tend toward similar values of scalar charge as they approach their end points.
+By comparing the branches, we found indications of which Maxwell coupling may stabilize solutions with quadratic coupling as the Ricci coupling and self-interactions did.
 Finally, we found that our model permits overcharged black hole solutions (with charge $Q$ exceeding the extremal value $M$).
@@
 One first step should derive the black hole solutions for $\eta<0$ ($GB^{-}$).
-The next one will be to study the stability of these solutions (radial and angular), the work of [102] make it simple to do the analysis.
+The next one will be to study the stability of these solutions (radial and angular), where the work of Ref [106] will make it simple to do the analysis.
 Future work can examine different coupling functions in our model.
@@
 Our new model with both curvature and matter couplings broadens the domain of spontaneous black hole scalarization and reveals a much richer structure of solutions.
-We have shown that scalarized black holes in this theory can exist under more flexible conditions, they appear to be more stable, possess higher entropy, and even achieve charge-to-mass ratios beyond the traditional extremal limit.
+We have shown that scalarized black holes in this theory can exist under more flexible conditions, where they appear to be more stable, possess higher entropy, and even achieve charge-to-mass ratios beyond the traditional extremal limit.
 These findings deepen our understanding of scalarization phenomena by introducing a new type of scalarization induced by two different sources and set the stage for further investigations into the stability and astrophysical implications of scalarized black holes.
@@
 $\quad-2rf_{,\phi}(\phi)\left(Q^{4}+2r^{3}(1-4N)g_{,\phi}(\phi)\phi^{\prime}Q^{2}+2r^{2}g(\phi)\left(-r^{2}\phi^{\prime 2}+N\left(\left(5r^{2}-4(N-1)f_{,\phi\phi}(\phi)\right)\phi^{\prime 2}+3\right)-3\right)Q^{2}\right)$
-$\quad-2rf_{,\phi}(\phi)\left(r^{4}g(\phi)^{2}\left(2r^{2}\phi^{\prime 2}\right.\right.$
+[…]
 $\quad\quad\left.\left.+N\left(-4(N-1)\left(r^{2}(3N-1)\phi^{\prime 2}+3(N-1)\right)f_{,\phi\phi}(\phi)\phi^{\prime 2}-2\left(r^{2}\phi^{\prime 2}+3\right)-N\left(r^{4}\phi^{\prime 4}+24r^{2}\phi^{\prime 2}-3\right)\right)+3\right)\right)$
 $\quad-r^{5}\left(g_{,\phi}(\phi)Q^{2}+2rg(\phi)\phi^{\prime}\left(r^{2}g(\phi)\left(2(N-1)Nf_{,\phi\phi}(\phi)\phi^{\prime 2}+N+1\right)-Q^{2}\right)\right),$
-$Z$ $=-2r^{2}g(\phi)N\left(8Q^{2}(N-1)f_{,\phi}(\phi)^{2}\left((3N+1)f_{,\phi}(\phi)\phi^{\prime}-r\right)+r^{2}g(\phi)\left(r^{5}+2f_{,\phi}(\phi)\left((1-7N)\phi^{\prime}r^{4}\right.\right.\right.$
+[…]
 $\quad\left.\left.\left.+6f_{,\phi}(\phi)\left(2r^{2}N(3N-1)\phi^{\prime 2}-(N-1)^{2}\right)r+4Nf_{,\phi}(\phi)^{2}\phi^{\prime}\left(6(N-1)^{2}+r^{2}(3N(2-5N)+1)\phi^{\prime 2}\right)\right)\right)\right).$